### v1

# Linearization of Monge-Ampère equations and data science applications

Abstract
Optimal transport has proven to have numerous applications in data science. In some of those, differentiating the transport map with respect to the densities in the Fréchet sense is required. In this work, we prove that for a compact, uniformly convex set $\Omega\subset\mathbb{R}^{d}$ and a $\mathcal{C}^{1}$ curve of densities $\{p_{t}\}\subset\mathcal{C}^{1,\alpha}(\Omega)$ that are uniformly bounded away from zero and infinity, the curve of transport maps $\{\nabla\phi_{t}\}\subset\mathcal{C}^{1,\alpha}(\Omega)$, pushing $Q$ forward to $P_{t}$ (where $P_{t}$ admits density $p_{t}$ and $Q$ is sufficiently regular), is also $\mathcal{C}^{1}$. The proof is based on applying the implicit function theorem to the Monge-Ampère equation with natural boundary conditions. To achieve this, we establish the existence and uniqueness of solutions to the linearized Monge-Ampère equation, an elliptic partial differential equation characterized by strictly oblique boundary conditions and a null zero-order term. We emphasize two significant applications of this result. Firstly, it enables non-parametric inference on the Monge map, paving the way for the construction of confidence bands for the transport map. Secondly, it establishes the regularity of the transport-based quantile regression function with respect to the covariates.
Keywords Monge-Ampère, Quantile regression, Oblique Boundary Conditions, Optimal Transport
AMS 2020 Subject Classification 35J96; 49Q22; 62G08
1 Introduction
The optimal transport problem has increasingly become a common research topic in data science, frequently appearing in both statistical journals (see e.g., [7, 16, 29]) and machine learning conferences (see e.g., [26, 1]). In all these applications, the true probabilistic model, i.e., the underilying probabilty measures $P$ and $Q$, are not observed. The practitioner have only access to a sample $X_{1},\dots,X_{n}$ of $P$ and $Y_{1},\dots,Y_{n}$ of $Q$. Let $p$ and $\Omega^{\prime}$ denote the density and support of $P$, respectively, and let $q$ and $\Omega$ denote the density and support of $Q$, respectively. Hence, the so-called Monge map $T_{Q\to P}$, which is the solution to the optimization problem
[…]
$a_{n}(T_{Q_{n}\to P_{n}}-T_{Q\to P})\implies G,$
where $\implies$ denotes weak convergence (in the probabilistic sense), $\{a_{n}\}_{n\in\mathbb{N}}$ is an increasing sequence of positive numbers, and $G$ is a non-trivial Gaussian process. However, $T_{Q_{n}\to P_{n}}$ suffers from the curse of dimensionality. That is, the rate of convergence $a_{n}$ at which $T_{Q_{n}\to P_{n}}$ approaches $T_{Q\to P}$ becomes progressively slower as the dimension increases [20]. In fact, the literature offers satisfactory results only in one-dimensional settings [15].
In nonparametric statistics, for example, in density estimation [41] or regression [27], this is a common problem. This is generally approached by making a priori assumptions about the smoothness of the population function—whether it be the regression function or the density $p$ of $P$—that is to be estimated, and by designing estimators that match this level of smoothness (see [23]). The smoothness of the target functions plays a crucial role in estimation problems. Broadly speaking, it can be understood that according to empirical process theory [40], smoother functions tend to yield faster rates of convergence (Theorem 2.7.1, Idem).
For this reason, several statisticians (see e.g., [35, 12]) have proposed estimating the population Monge map by computing the Monge map $T_{h_{n}}$ between the empirical measures convoluted with a kernel $K_{h_{n}}$ density estimator [^1] Recall from [41, Section 4.2], for all $z\in\mathbb{R}^{d}$, the kernel $K_{h_{n}}(z-\cdot):x\mapsto K_{h_{n}}(z-x)$ is non-negative and integrates to 1. having bandwidth $h_{n}$. To be more specific, the map $T_{h_{n}}=\nabla\phi_{h_{n}}$ solves the Monge-Ampère equation
$\det(D^{2}\phi_{h_{n}})$ $=\frac{K_{h_{n}}*Q_{n}}{[K_{h_{n}}*P_{n}](\nabla\phi_{h_{n}})}\quad\rm in\,\,{\rm supp}(K_{h_{n}}*Q_{n}),$ (2)
$\nabla\phi_{h_{n}}({\rm supp}(K_{h_{n}}*Q_{n}))$ $={\rm supp}(K_{h_{n}}*P_{n}),$
where
$[K_{h_{n}}*P_{n}](x)=\frac{1}{n}\sum_{i=1}^{n}K_{h_{n}}(X_{i}-x)\quad\text{and}\quad[K_{h_{n}}*Q_{n}](y)=\frac{1}{n}\sum_{i=1}^{n}K_{h_{n}}(Y_{i}-y).$
In the nomenclature of empirical processes, $T_{h_{n}}$ can be seen as the solution of a Z-estimation problem. One way to obtain asymptotic confidence bands (and rates of convergence) of $T_{h_{n}}-T$ is by linearizing (2). Specifically, this involves applying an implicit function theorem to express $T_{P\to Q}\in\mathcal{C}^{1,\alpha}(\Omega)$ as a differentiable functional of the densities $p\in\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ and $q\in\mathcal{C}^{1,\alpha}(\overline{\Omega})$ of $P$ and $Q$, respectively. Once this is established, it will be possible to obtain expressions of the type
$T_{h_{n}}=T_{Q\to P}+L((K_{h_{n}}*P_{n}-p),(K_{h_{n}}*Q_{n}-q))\\ +o_{P}(\|K_{h_{n}}*P_{n}-p\|_{\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})}+\|K_{h_{n}}*Q_{n}-q\|_{\mathcal{C}^{1,\alpha}(\overline{\Omega})}),$ (3)
where is a bounded linear operator. This readily provides bounds of type
$\|T_{h_{n}}-T_{Q\to P}\|_{\mathcal{C}^{1,\alpha}(\Omega)}\leq C\left(\|K_{h_{n}}*P_{n}-p\|_{\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})}+\|K_{h_{n}}*Q_{n}-q\|_{\mathcal{C}^{1,\alpha}(\overline{\Omega})}\right).$ (4)
and if $a_{n}(K_{h_{n}}*P_{n}-p)$ and $a_{n}(K_{h_{n}}*Q_{n}-q)$ have weak limits in $\mathcal{C}^{\alpha}(\Omega^{\prime})$ and $\mathcal{C}^{\alpha}(\Omega)$, respectively, one gets the asymptotic distribution of $a_{n}(T_{h_{n}}-T_{Q\to P})$ via the functional delta method. Interested reader might observe that, when $Q$ is fixed, bounds of type (4) can be readily deduced from the results of this paper. To obtain confidence intervals for $T_{h_{n}}$, a precise study of kernel density estimators is required, which falls outside the scope of this work. We refer interested readers to the monograph [23] for further discussion on smooth density estimation.
[35, 34] employed a similar strategy to obtain point-wise confidence bands for $T_{h_{n}}$ for probability measures on the flat torus. It can be noted that the asymptotic expansion (3) corresponds to [34, Theorem 2], while the bound (4) corresponds to (Proposition 12, Idem). Their restriction to the flat torus was a consequence of the difficulty in linearizing the Monge-Ampère equation in the Euclidean case due to boundary conditions. Optimal transport on the flat torus, despite its applications in areas such as structural biology [24, 25], has a narrower range of applications compared to the Euclidean case. This limits its broader utilization and effectiveness.
We now consider another example where linearizing the Monge-Ampère equation is crucial: quantile regression. Introduced by Koenker and Bassett [33], quantile regression explains the influence of a random variable $X$ (the covariates) on another random variable $Y$ (the output or variable of interest) through the conditional quantile function $Q_{Y|X}(\cdot,\cdot)$ of $Y$ given $X$. Unlike other models, it provides a complete view of the conditional distribution, capturing both trends and heteroscedasticity. However, generalizing $Q_{Y|X}$ to multivariate outputs is challenging due to the lack of canonical ordering in $\mathbb{R}^{d}$, which does not allow a direct definition of multivariate (conditional) quantiles (see [28]). Recently, [29, 10] have made significant progress in defining multivariate quantiles, which has proven useful in various applications [21, 38, 2, 16]. Such an approach defines the transport-based quantile function $\mathbb{Q}$ of a probability measure $P$ with density $p$ (i.e., $\mathrm{d}P=\mathbf{1}_{\Omega^{\prime}}p\,\mathrm{d}x$) as the unique gradient of a convex function $\varphi$ that pushes forward the reference measure $Q$ with density $q$ (supported on a convex set $\Omega$) to $P$. Specifically, $\varphi$ is a convex solution to the Monge-Ampère equation:
$\det(D^{2}\varphi)=\frac{q}{p(\nabla\varphi)}\quad\text{in}\ \Omega,\quad\nabla\varphi(\Omega)=\Omega^{\prime}.$
The inverse, namely $\mathbb{F}$, of $\mathbb{Q}$ is called the transport-based distribution function [^2] In [29], the term center-outward distribution function was introduced for $\mathbb{F}$, specifically when the reference measure $Q$ is chosen to be the spherical uniform distribution. , which is the gradient of $\phi=\varphi^{*}:=\sup_{x\in\mathbb{R}^{d}}\{\left\langle x,\cdot\right\rangle-\varphi(x)\}$ (the convex conjugate $\varphi$).
Regarding the quantile regression problem, when we consider two random vectors $(X,Y)\in\mathbb{R}^{m+d}$ with probability law $P_{X,Y}$, the quantile regression function of $X$ on $Y$ is the unique Borel function $\mathbb{Q}_{Y|X}:\Omega^{\prime}\times\mathbb{R}^{m}\to\mathbb{R}^{d}$ such that for $P_{X}$-almost every $x$, the function $u\mapsto\nabla_{u}\varphi(u,x):=\mathbb{Q}_{Y|X}(u,x)$ is the quantile function of $P_{Y|X=x}$ —the conditional probability measure of $Y$ given $X=x$— having density $p_{Y|X}(\cdot,x)$ (see [7, 16]). Therefore $\varphi(u,x)$ solves the conditional Monge-Ampère equation
$\det(D^{2}_{u}\varphi(u,x))=\frac{q(u)}{p_{Y|X}(\nabla_{u}\varphi(u,x),x)}\quad\text{for}\,\,u\in\ \Omega,\quad\nabla\varphi(\Omega\times\{x\})=\Omega_{x}:={\rm supp}(P_{Y|X}(\cdot,x)).$
Applications of this conditional Monge map (or quantile regression function), apart from those well-known of quantile regression (see [16] and references therein), range from Bayesian inverse problems (see [31, 32]) to transformers in Machine learning (see [8]).
The regularity of $\varphi(u,x)$ with respect to the variable $u$ is well-understood and follows as a consequence of Caffarelli’s regularity theory (see [4, 5, 18, 19, 14]). However, little is known about the regularity of $\varphi(u,x)$ with respect to the covariates (the variable $x$). As highlighted in [31], addressing this issue is crucial for advancing the practical applications and theoretical understanding of the conditional Monge maps. In this work, we provide sufficient conditions to obtain this regularity.
We observe that from a broader perspective, both the regularity of the quantile regression function and the derivation of confidence intervals for $T_{h_{n}}$ (the solution of (2)) can be framed as follows. Let $I\subset\mathbb{R}$ be an interval, and let $\{P_{t}\}_{t\in I}$ denote a family of probability measures with densities $\{p_{t}\}_{t\in I}$ and supports $\{\Omega_{t}\}_{t\in I}$. The objective is to identify sufficient conditions on $\{P_{t}\}_{t\in I}$ and $Q$ such that the curve $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\Omega)$ solving
$\det(D^{2}\phi_{t})$ $=\frac{q}{p_{t}(\nabla\phi_{t})}\quad\text{in }\Omega,$ (5)
$\nabla\phi_{t}(\Omega)$ $=\Omega_{t}$
is $\mathcal{C}^{1}$. We state here a simplified version of our main result (see Theorem 3.2), where we assume that the supports of $\{P_{t}\}_{t\in I}$ do not vary with $t$.
Theorem 1.1.
Let $\Omega$ and $\Omega^{\prime}$ be $\mathcal{C}^{2}$ uniformly convex domains. Let $\log(p_{t})\in\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ with $0<\gamma\leq 1$ and let $\log(q)\in\mathcal{C}^{\alpha}(\overline{\Omega})$ with $0<\alpha<\gamma$ such that $11=\int_{\Omega}q=\int_{\Omega^{\prime}}p_{t}=$ 1 for all $t\in I$. If the curve $\{\log(p_{t})\}_{t\in I}\subset\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ is $\mathcal{C}^{1}$, then the curve $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\overline{\Omega})$, where $\phi_{t}$ solves (5) for $\Omega_{t}=\Omega^{\prime}$, is also $\mathcal{C}^{1}$. Moreover, the time derivative of $I\ni t\mapsto\phi_{t}\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$, denoted as $\partial_{t}\phi_{t}$, is the unique solution (up to an additive shift) of the linearized Monge-Ampère equation
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=-\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}\quad\rm in\,\,\Omega,$ (6)
$\langle\nu_{\Omega^{\prime}}(\nabla\phi_{t}),\nabla\xi\rangle$ $=0\quad\rm on\,\,\partial\Omega,$
where $\nu_{\Omega^{\prime}}$ denotes the unit outer normal vector-field to $\partial\Omega^{\prime}$.
From Theorem 1.1, we obtain as a corollary the aforementioned regularity of the quantile regression function. As we mentioned earlier, this can be applied to obtain asymptotic confidence intervals for the transport map. However, we will leave this topic for future work.
Corollary 1.2.
Let $\Omega$, $\Omega^{\prime}$ and the reference measure $Q={\bf 1}_{\Omega}qdx$ be as in Theorem 1.1. Let $(X,Y)\in\mathbb{R}^{m+d}$ be a pair of random variables with probability law $P_{X,Y}=P_{Y|X}P_{X}$. Assume that
  • for each $x\in{\rm int}({\rm supp}(P_{X}))$, the conditional probability measure has support $\Omega^{\prime}$ and density $p_{Y|X}(\cdot,x)\in\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ (i.e., $P_{Y|X=x}={\bf 1}_{\Omega^{\prime}}p_{Y|X}(\cdot,x)dy$) and
  • the mapping ${\rm int}({\rm supp}(P_{X}))\ni x\mapsto\log(p_{Y|X}(\cdot,x))\in\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ is $\mathcal{C}^{1}$.
Then the function ${\rm int}({\rm supp}(P_{X}))\ni x\mapsto\mathbb{Q}_{Y|X}(\cdot,x)\in\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ is $\mathcal{C}^{1}$. As a consequence, $\nabla_{x}\mathbb{Q}_{Y|X}\in\mathcal{C}(\overline{\Omega^{\prime}}\times{\rm supp}(P_{X}))$.
The proof of our main result is based on the application of the implicit function theorem to the functional
$\Gamma(t,\phi)=\left(\begin{array}[]{c}\Gamma_{t}^{(1)}(\phi)\\ \Gamma_{t}^{(2)}(\phi)\end{array}\right)=\left(\begin{array}[]{c}\log(\det(D^{2}\phi))-\log(q)+\log(p_{t}(\nabla\phi))\\ h(\nabla\phi)\end{array}\right),$
where $h$ is a uniformly convex defining function of the support set $\Omega^{\prime}$. The functional $\Gamma(t,\cdot)$ is defined over the set of strongly convex $\mathcal{C}^{2,\alpha}(\Omega)$ functions. The range of $\Gamma$ is $\mathcal{C}^{0,\alpha}(\Omega)\times\mathcal{C}^{1,\alpha}(\partial\Omega)$ functions. To apply the aforementioned implicit function theorem one need to show the invertibility of the Fréchet derivative of $\Gamma$ w.r.t. $\phi$. The proof of the existence of continuous Fréchet derivatives follows the essential steps outlined in [39, 18], with the exception of a few subtleties that ensure the functional is always well-defined in their respective spaces (see Lemma 4.5). The invertibility involves studying the following second-order elliptic differential equation with strictly oblique boundary conditions:
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=f\quad\rm in\,\,\Omega,$
$\langle\nu_{\Omega^{\prime}}(\nabla\phi_{t}),\nabla\xi\rangle$ $=g\quad\rm on\,\,\partial\Omega.$
We will show in Theorem 4.6 that the precedent elliptic operator defines a bijection between the spaces
$\mathcal{X}_{t}=\left\{\xi\in\mathcal{C}^{2,\alpha}(\overline{\Omega}):\int_{\partial\Omega^{\prime}}p_{t}\xi(\nabla\phi_{t}^{*})=0\right\}$
and
$\mathcal{Y}_{t}:=\left\{(f,g)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega):\int_{\Omega}qf=\int_{\partial\Omega^{\prime}}p_{t}g(\nabla\phi_{t}^{*})\right\}.$
Here, we encounter a minor obstacle; the range of $\Gamma$ exceeds the set $\mathcal{Y}_{t}$. Therefore, for each fixed $t$, we redefine $\Gamma_{s}$ as
$\Gamma^{(t)}(s,\phi):=\begin{pmatrix}\Gamma_{s}^{(1)}(\phi)-\int_{\Omega}q\Gamma_{s}^{(1)}(\phi)+\int_{\partial\Omega_{t}}p_{t}[\Gamma_{s}^{(2)}(\phi)](\nabla\phi_{t}^{*})\\ \Gamma_{s}^{(2)}(\phi)\end{pmatrix}$ (7)
to reduce its range (see Lemma 4.11), allowing us to apply the implicit function theorem to $\Gamma^{(t)}$ around the point $(t,\phi_{t})$ and conclude the proof.
The remainder of the paper is organized as follows: Section 2 provides and recalls part of the notation used throughout the work for the reader’s convenience. In Section 3, the main result and its assumptions are stated. Section 4 is dedicated to the proof of the main result.
2 Notation
For any Borel probability measure $P$ over $\mathbb{R}^{d}$ (i.e., $P\in\mathcal{P}(\mathbb{R}^{d})$), let $\mathrm{supp}(P)$ denote its topological support, which is defined to be
$\mathrm{supp}(P):=\mathbb{R}^{d}\setminus\bigcup\{U\subseteq\mathbb{R}^{d}:U\text{ is open },P(U)=0\}.$
[…]
Let $U\subseteq\mathbb{R}^{d}$ be bounded, $k\in\mathbb{N},\alpha\in(0,1]$, we denote by $\mathcal{C}^{k,\alpha}(U)$ the space consisting of functions whose $k$-th order partial derivatives are uniformly Hölder continuous with exponent $\alpha$. Let $f\in\mathcal{C}^{1}(U)$, $e\in\mathbb{R}^{d}$, we denote by $\partial_{e}u$ the directional derivative along the direction $e$ and denote by $\nabla f:=(\partial_{e_{1}}f,\dots,\partial_{e_{d}}f)^{\intercal}$, where $\{e_{j}\}_{j=1}^{d}$ denotes the canonical basis of $\mathbb{R}^{d}$. Hessian matrix of a function $f\in\mathcal{C}^{2}(U)$ is denoted as $D^{2}f$. Let $F:\mathbb{R}^{d}\to\mathbb{R}^{d}$, we denote by $\,\mathrm{div}(F)$ the divergence of $F$. Denote by $\mathbb{R}_{+}^{d}$ the half-space $\{x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}:x_{d}>0\}$. We call domain every open, bounded, connected, and non-empty subset $\Omega$ of $\mathbb{R}^{d}$. Moreover, $\Omega$ is said to be a $\mathcal{C}^{k,\alpha}$ domain $(k\in\mathbb{N},\alpha\in[0,1])$ if for every $p\in\partial\Omega$, the boundary of $\Omega$, there exists a neighborhood $B=B(x_{0})$ of $x_{0}$ in $\mathbb{R}^{d}$ and a diffeomorphism $\Psi:B\rightarrow D\subset\mathbb{R}^{d}$ such that (i) $\Psi(B\cap\Omega)\subset\mathbb{R}_{+}^{d}$; (ii) $\Psi(B\cap\partial\Omega)\subset\partial\mathbb{R}_{+}^{d}$; (iii) $\Psi\in\mathcal{C}^{k,\alpha}(B),\Psi^{-1}\in\mathcal{C}^{k,\alpha}(D)$.
We denote by ${\rm tr}(M)$ the trace of a matrix $M$. Given two symmetric matrices $A,B$ with dimension $d\times d$ over $\mathbb{R}$, we say that $A\leq B$ if $B-A$ is positive semi-definite.
3 Main result
In this section, we state the main result of this work and introduce its assumptions. We consider a curve $\{P_{t}\}_{t\in I}$ of probability measures with both smooth (and uniformly convex) supports and densities. Recall that a curve $\{b_{t}\}_{t\in I}$ in a Banach space $(B,\|\cdot\|_{B})$ is $\mathcal{C}^{1}$ if there exists a continuous function $\partial_{t}b_{t}:I\to B$ such that the limit $\frac{1}{h}(b_{t+h}-b_{t})\xrightarrow{h\to 0}\partial_{t}b_{t}$ holds in norm for all $t\in{\rm int}(I)$.
As mentioned in the introduction, our main result is more general than the previously stated Theorem 1.1, as it accommodates time-dependent supports. Due to this dependency, the curve $\{p_{t}\}_{t\in I}$ is not well-defined in $\mathcal{C}^{1,\gamma}(\Omega^{\prime})$. Initially, we will assume that $\{p_{t}\}_{t\in I}$ has a $\mathcal{C}^{1,\gamma}$ extension. Later, in Lemma 3.4, we will provide sufficient conditions for this extension.
First, we define our notion of a $\mathcal{C}^{1}$ curve of probability measures in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ for a given set $\Omega^{\prime}$.
Definition 3.1.
A curve $\{P_{t}\}_{t\in I}$ of probability measures over a domain $\Omega^{\prime}$ is said to be $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ if the following conditions are satisfied:
  1. 1.
    There exists $\kappa\in(0,+\infty)$ and a $\mathcal{C}^{1}$ curve of convex function $\{h_{t}\}_{t\in I}$ in $\mathcal{C}^{2,\alpha}(\overline{\Omega^{\prime}})$ such that
    $\Omega_{t}:={\rm int}(\mathrm{supp}(P_{t}))=\left\{y\in\mathbb{R}^{d}:h_{t}(y)<0\right\},$
    with $\|\nabla h_{t}\|=1$ on $\partial\Omega_{t}$, $\bigcup_{t\in I}\Omega_{t}+\frac{1}{\kappa}\mathbb{B}=\left\{y:{\rm dist}\left(y,\bigcup_{t\in I}\Omega_{t}\right)<\frac{1}{\kappa}\right\}\subset\Omega^{\prime}$, and
    $\frac{1}{\kappa}{\rm I}_{d}\leq D^{2}h_{t}\leq\kappa{\rm I}_{d}\quad\text{in }\Omega^{\prime}.$
  2. 2.
    There exists a $\mathcal{C}^{1}$ curve $\{\log(p_{t})\}_{t\in I}$ in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ such that $P_{t}=p_{t}{\bf 1}_{\Omega_{t}}$.
The function $h_{t}$ is commonly known as the convex defining function of $\mathrm{supp}(P_{t})$. It is noteworthy that $\Omega_{t}=\left\{y\in\mathbb{R}^{d}:h_{P}(y)<0\right\}$ represents an open, bounded, and uniformly $\mathcal{C}^{2,\alpha}$ convex domain. Conversely, given a uniformly convex set $\Omega_{t}$, one can construct the function $h_{t}$ to be smooth and uniformly convex, approximating $-\text{dist}(\cdot,\partial\Omega_{t})+\frac{1}{2}\text{dist}(\cdot,\partial\Omega_{t})^{2}$ near $\partial\Omega_{t}$. For example, if $\Omega_{t}=\mathbb{B}(0,1)=\{x\in\mathbb{R}^{d}:\|x\|=1\}$, one might choose $h_{P}(y):=\frac{1}{2}(\|y\|^{2}-1)$ (see e.g., [18, pp 40]).
Theorem 3.2.
Let $\Omega$ be a $\mathcal{C}^{2}$ uniformly convex domain and let $\Omega^{\prime}\subset\mathbb{R}^{d}$ be a domain. If the curve $\{P_{t}\}_{t\in I}$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ with $0<\gamma\leq 1$ (in the sense of Definition 3.1) and let $\log(q)\in\mathcal{C}^{\alpha}(\overline{\Omega})$ with $0<\alpha<\gamma$, then the curve $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\overline{\Omega})$, where $\phi_{t}$ solves (5), is also $\mathcal{C}^{1}$. Moreover, the time derivative of $I\ni t\mapsto\phi_{t}\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$, denoted as $\partial_{t}\phi_{t}$, is the unique solution (up to an additive shift) of the linearized Monge-Ampère equation
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=-\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}\quad\rm in\,\,\Omega,$ (8)
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=-\partial_{t}h_{t}(\nabla\phi_{t})\quad\rm on\,\,\partial\Omega.$
Remark 3.3.
Recent advances in the Monge-Ampère equation [9] suggest that the assumptions on $\Omega$ and $\{\Omega_{t}\}_{t\in I}$ could be relaxed to being convex and $\mathcal{C}^{1,1}$. However, to avoid additional technical complications in the proof, we adhere to the conditions assumed in [6, 39].
As we mentioned earlier, we are implicitly assuming that the densities $p_{t}$, typically known only on $\Omega_{t}$, admit an extension to $\Omega^{\prime}$. In the following result, we provide a sufficient condition for this extension to exist, thereby expanding the range of applications of Theorem 3.2. We establish the proof of the following result below, while the proof of Theorem 3.2 is postponed to the next section.
Lemma 3.4.
Assume that $I$ is open. Let $\{h_{t},\Omega_{t}\}_{t\in I}$ and $\Omega^{\prime}$ be defined as in Definition 3.1, let $h:(t,x)\mapsto h(t,x)=h_{t}(x)$ be $\mathcal{C}^{1,\alpha}$ over $I\times\Omega^{\prime}$, and let $\log(p):(t,x)\mapsto\log(p(t,x))=\log(p_{t}(x))$ be a $\mathcal{C}^{1,\alpha}$ function over the set $\mathcal{D}:=\{t,\mathbf{x}):\ t\in I,\ \mathbf{x}\in\overline{\Omega}_{t}\}.$ Then the curve $\{P_{t}\}_{t\in I}$ with $P_{t}=p_{t}{\bf 1}_{\Omega_{t}}$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ in the sense of Definition 3.1.
Proof.
We want to extend $\log p$ from ${\rm int}(\mathcal{D}):=$${\rm int}\left(\{(t,x):\ t\in I,\ x\in\Omega_{t}\}\right)$ to $I\times\Omega^{\prime}$ using [22, Lemma 6.37]. Since $I$ is open, we may assume without loss of generality that $I=\mathbb{R}$. Then, for every $(t_{0},x_{0})\in\partial D$, it suffices to find a neighborhood $B(t_{0},x_{0})$ of $(t_{0},x_{0})$ and a $\mathcal{C}^{1,\alpha}$ diffeomorphism $\Psi:B(t_{0},x_{0})\to D\subset\mathbb{R}^{d+1}$ such that $\Psi(B(t_{0},x_{0})\cap\mathcal{D})\subset\mathbb{R}^{d+1}_{+}$, $\Psi(B(t_{0},x_{0})\cap\partial\mathcal{D})\subset\partial\mathbb{R}^{d+1}_{+}$, and $\Psi\in\mathcal{C}^{1,\alpha}(B(t_{0},x_{0})),\Psi^{-1}\in\mathcal{C}^{1,\alpha}(D)$. A candidate map is
$\Psi:\mathbb{R}^{d+1}\ni(t,x_{1},\cdots,x_{d-1},x_{d})\mapsto(t,x_{1},\cdots,x_{d-1},-h_{t}(x))\in\mathbb{R}^{d+1}.$
Since $h_{t}$ is a convex defining function of $\Omega_{t}$, the first two requirements are easily satisfied, and $\Psi$ is $\mathcal{C}^{1,\alpha}$. We now show that $\Psi$ is one-to-one in a small neighbourhood of $(t_{0},x_{0})$. As $\{x\in\Omega:h_{t}(x)<0\}$ is open and $h_{t}$ is uniformly convex, one may choose $\delta$ small enough such that $\|\nabla h_{t}(x)\|_{2}\neq 0$ for all $(t,x)\in B:=\mathbb{B}((t_{0},x_{0}),\delta)$. We proceed with the proof by contradiction. Suppose there exists $(t,x)=(t,x_{1},\cdots,x_{d})\neq(s,y)=(s,y_{1},\cdots,y_{d})\in B$ such that $\Psi(t,x)=\Psi(s,y)$, then
$t=s,x_{i}=y_{i},\quad i=1,\cdots,d-1,$
and $h_{t}(x)=h_{t}(y)$. Define $\tilde{h}:z\mapsto h_{t}(x_{z})$, where $x_{z}:=(x_{1},\cdots,x_{d-1},z)$, then $\tilde{h}$ is also $\mathcal{C}^{1,\alpha}$ uniformly convex. Moreover, as $|\tilde{h}^{\prime}(z)|\neq 0$ over $\mathcal{Z}:=\{z:(t,x_{z})\in B\}$, $\tilde{h}^{\prime}(z)=\left\langle\nabla h_{t}(x_{z}),e_{d}\right\rangle$ is either uniformly positive or uniformly negative over $\mathcal{Z}$ by the monotonicity of the gradient of a uniformly convex function. Therefore, the fundamental theorem of calculus yields that
[…]
which leads to a contradiction. ∎
4 Proof of the Main Result and solvability of linearized equation
In this section, we provide the proof of Theorem 3.2. We recall that for each $t$, the function $\phi_{t}:\Omega\to\mathbb{R}$ is strictly convex and solves the boundary problem (5). Following [39], we rewrite (5) in such a way that a linearization is viable:
$\det(D^{2}\phi_{t})$ $=\frac{q}{p_{t}(\nabla\phi_{t})}\quad{\rm in}\,\,\Omega,$ (9)
$h_{t}(\nabla\phi_{t})$ $=0\quad{\rm in}\,\,\partial\Omega.$
As indicated previously, the strategy is to differentiate (9) and apply the implicit function theorem [13, Theorem 15.1] over the functional
$\Gamma:I\times\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})$ $\to\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$
$\left(\begin{array}[]{c}t\\ \phi\end{array}\right)$ $\mapsto\left(\begin{array}[]{c}\Gamma_{t}^{(1)}(\phi)\\ \Gamma_{t}^{(2)}(\phi)\end{array}\right):=\left(\begin{array}[]{c}\log(\det(D^{2}\phi))-\log(q)+\log(p_{t}(\nabla\phi))\\ h_{t}(\nabla\phi)\end{array}\right),$ (16)
where for $\beta\geq 0$,
$\mathcal{C}^{k,\alpha}_{>\beta}(\bar{\Omega}):=\{f\in\mathcal{C}^{k,\alpha}(\overline{\Omega}):D^{2}f>\beta I\}.$
We observe that the terms $\log(p_{t}(\nabla\phi))$ and $h_{t}(\nabla\phi)$ appear in the definition of $\Gamma$. Therefore, we must ensure that the composition of Hölder continuous functions with the same exponents remains Hölder continuous with exponent unchanged. We note that such a claim is untrue in general. For example, if $f,g\in\mathcal{C}^{0,\alpha}(\overline{\Omega})$, then $f\circ g\in\mathcal{C}^{0,\alpha^{2}}(\overline{\Omega})$ rather than being $\mathcal{C}^{0,\alpha}(\overline{\Omega})$. However, the claim is valid when either $f$ or $g$ has higher regularity as in the following lemma, which is an immediate result from [11, Theorem 4.3].
Lemma 4.1 (Compositions of Hölder Functions).
For any $\alpha,\beta\in(0,1],l\geq k\geq 0,l\geq 1$, there exists a constant $C=C(k,l,\alpha,\beta,\rm diam(\Omega))>0$ such that for all $f\in\mathcal{C}^{k,\alpha}(\Omega^{\prime})$ and $g\in\mathcal{C}^{l+1,\beta}(\overline{\Omega})$ such that $\nabla g(\overline{\Omega})\subset\Omega^{\prime}$, then $f\circ\nabla g\in\mathcal{C}^{k,\min(\alpha,\beta)}(\overline{\Omega})$ and
$\|f\circ\nabla g\|_{\mathcal{C}^{k,\min(\alpha,\beta)}(\overline{\Omega})}\leq C\left(1+\|g\|_{\mathcal{C}^{l+1,\beta}(\overline{\Omega})}^{\alpha}\right)\|f\|_{\mathcal{C}^{k,\alpha}(\Omega^{\prime})}.$
As a direct consequence of Lemma 4.1, we establish that $\Gamma$ is well-defined. We conclude this subsection by presenting a set of properties related to Hölder spaces, which will be instrumental in the subsequent subsection for proving that the Fréchet derivative of $\Gamma$ with respect to $\phi$ is well-defined. The following lemma, adapted from [11, Proposition 6.1,6.2], states the stability of Hölder functions under composition, and justifies the condition in Theorem 3.2, where we assume that $p_{t}$ is $\mathcal{C}^{1,\gamma}$ rather than merely $\mathcal{C}^{1,\alpha}$.
Lemma 4.2.
Under the same setting as in Lemma 4.1, the map
$\mathcal{C}^{k,\alpha}(\Omega^{\prime})\ni f\mapsto f\circ\nabla g\in\mathcal{C}^{k,\min{(\alpha,\beta)}}(\overline{\Omega})$
is linear and continuous. In addition, if $\alpha>\beta$, then the map
$\mathcal{C}^{l+1,\beta}(\overline{\Omega})\ni g\mapsto f\circ\nabla g\in\mathcal{C}^{k,\beta}(\overline{\Omega})$
is also continuous. More precisely, there exists $\delta,\rho,M>0$ such that if $\left\lVert g^{\prime}-g\right\rVert_{\mathcal{C}^{l+1,\beta}(\overline{\Omega})}<\delta$, one has that
$\|f\circ\nabla g-f\circ\nabla g^{\prime}\|_{\mathcal{C}^{k,\beta}(\overline{\Omega})}\leq M\|f\|_{\mathcal{C}^{k,\alpha}(\Omega^{\prime})}\|g-g^{\prime}\|_{\mathcal{C}^{l+1,\beta}(\overline{\Omega})}^{\rho}.$
The next result is standard and can be found in [22, Section 4.1].
Lemma 4.3 (Product of Hölder Functions).
For any $\alpha,\beta\in(0,1],l\geq k\geq 0$, there exists a constant $C=C(k,l,\alpha,\beta,\rm diam(\Omega))>0$ such that for all $f\in\mathcal{C}^{k,\alpha}(\overline{\Omega}),g\in\mathcal{C}^{l,\beta}(\overline{\Omega})$,
$\left\lVert fg\right\rVert_{\mathcal{C}^{l,\min(\alpha,\beta)}(\overline{\Omega})}\leq C\left\lVert f\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}\left\lVert g\right\rVert_{\mathcal{C}^{l,\beta}(\overline{\Omega})}.$
As a consequence of Lemma 4.1 and Lemma 4.3, we immediately arrive at the following result as the function $1/g$ may be viewed as the composition of $x\mapsto 1/x$ and $g$, and $x\mapsto 1/x$ is $\mathcal{C}^{\infty}$ on $[\lambda,\infty)$.
Lemma 4.4.
Let $f,g\in\mathcal{C}^{k,\alpha}(\overline{\Omega})$ and $g\geq\lambda$ for some constant $\lambda>0$, then $f/g\in\mathcal{C}^{k,\alpha}(\overline{\Omega})$ and there exists $C=C(k,\alpha,\lambda,\rm diam(\Omega))$ such that
$\left\lVert f/g\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}\leq C\left\lVert f\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}\left\lVert g\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}.$
4.1 Fréchet derivative of the functional
Now we show that the map $\Gamma$ is $\mathcal{C}^{1}$. Recall that a mapping $F:\mathcal{U}\to Y$, were $(X,\|\cdot\|_{X})$ and $(Y,\|\cdot\|_{Y})$ are Banach spaces and $\mathcal{U}\subset X$ is open, is said to be Fréchet differentiable at $x\in\mathcal{U}$ if there exists a bounded linear functional $A(x):X\to Y$ such that
$\lim_{h\to 0}\frac{\left\lVert F(x+h)-F(x)-A(x)h\right\rVert_{Y}}{\left\lVert h\right\rVert_{X}}=0$
for every $h\in X$. We say that $F$ is Fréchet differentiable in $\mathcal{U}$ if $F$ is Fréchet differentiable at every point $x\in\mathcal{U}$. We say that $F$ is $\mathcal{C}^{1}$ in $\mathcal{U}$ if it is Fréchet differentiable in $\mathcal{U}$ and the mapping $A:X\to\mathcal{L}(X,Y)$, where $\mathcal{L}(X,Y)$ denotes the space of bounded linear functional from $X$ to $Y$, is continuous w.r.t. the norm topology. The proof follows the standard approach used in the continuity method (openness part) for the Monge-Ampère equation with Dirichlet boundary conditions (see [18]).
Lemma 4.5.
Under the assumptions of Theorem 3.2, the following holds:
  1. 1.
    $\Gamma_{t}(\phi_{t})=0$ for all $t\in I$.
  2. 2.
    For each $t\in I$ there exist an open set $I^{\prime}\times\mathcal{U}_{t}\subset I\times\mathcal{C}^{2,\alpha}_{>0}(\overline{\Omega})$ with $(t,\phi_{t})\in I^{\prime}\times\mathcal{U}_{t}$ such that $\Gamma$ is $\mathcal{C}^{1}$ in $I^{\prime}\times\mathcal{U}_{t}$.
  3. 3.
    For every $t\in I$, the Fréchet derivative of $\Gamma$ w.r.t. $\phi$ at $(t,\phi_{t})$ is the linear functional
    $D_{\phi}\Gamma_{t}(\phi_{t}):\mathcal{C}^{2,\alpha}(\overline{\Omega})$ $\to\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$
    $\xi$ $\mapsto\left(\begin{array}[]{c}{\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}\\ \langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle\end{array}\right),$
    which we call the linearized Monge-Ampère operator at $(t,\phi_{t})$.
  4. 4.
    For every $t\in I$, the Fréchet derivative of $\Gamma$ w.r.t. $t$ at $(t,\phi_{t})$ is the element
    $D_{t}\Gamma_{t}(\phi_{t})=\left(\begin{array}[]{c}\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}\\ \partial_{t}h_{t}(\nabla\phi_{t})\end{array}\right)$
    of $\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$.
Proof.
The first claim is straightforward by (5), $\nabla\phi_{t}(\partial\Omega)=\partial\Omega_{t}$, and $h_{t}(\partial\Omega_{t})=\{0\}$. We now prove the second and third claims together. Fix $t\in(0,1)$. The boundary regularity provided in [6, 39] yields the existence of a constant $C=C_{t}$ such that
$\|\phi_{t}\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq C\quad{\rm and}\quad\|\phi_{t}^{*}\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq C,$
which implies that $D^{2}\phi_{t}$ belongs to the set of (strictly) positive definite $d\times d$ real matrices over the field $\mathbb{R}$ (see also Lemma 4.8 below). Denote such set as $\mathcal{M}^{+}_{d\times d}$. Choose $\delta_{t}>0$ such that $D^{2}(\phi_{t}+\xi)\in\mathcal{M}^{+}_{d\times d}$ and $\nabla(\phi_{t}+\xi)(\overline{\Omega})\subset\Omega^{\prime}$ for all
$\xi\in\mathbb{B}_{\delta_{t}}:=\left\{f\in\mathcal{C}^{2,\alpha}(\overline{\Omega}):\|f\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq\delta_{t}\right\}.$
(Note that the latter has been possible due to the assumption $\bigcup_{t\in I}\Omega_{t}+\frac{1}{\kappa}\mathbb{B}\subset\Omega^{\prime}$.) Set $\mathcal{U}_{t}:=(\phi_{t}+\mathbb{B}_{\delta_{t}})\subseteq\mathcal{C}^{2,\alpha}_{>0}(\overline{\Omega})$. We show the second claim with the set $I^{\prime}\times\mathcal{U}_{t}=I\times\mathcal{U}_{t}$. Fix $(s,\phi)\in I\times\mathcal{U}_{t}$. It is well-known (see, e.g., [18, Section 3]) that
$\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\supset\mathcal{U}_{t}\ni\phi\mapsto\log\det(\nabla^{2}\phi)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})$
is $\mathcal{C}^{1}$ with directional derivative
$\frac{d}{d\varepsilon}\Big{|}_{\varepsilon=0}\log\det(A+\varepsilon B)=\rm tr(A^{-1}B).$
Since
$\Gamma(s,\phi)=\left(\begin{array}[]{c}-\log(q)+\log(p_{s}(\nabla\phi))\\ h_{s}(\nabla\phi)\end{array}\right)+\left(\begin{array}[]{c}\log(\det(D^{2}\phi))\\ 0\end{array}\right)$
and $\log(q)$ does not variate with $(s,\phi)$, we only need to prove the claims for
$(s,\phi)\mapsto(\log(p_{s}(\nabla\phi)),h_{s}(\nabla\phi)).$
Fix $\epsilon>0$. The chain rule [37, Proposition 3.6] and Lemma 4.1 imply the existence $\delta=\delta(\epsilon)>0$ such that the term $\phi\mapsto\log(p_{t}(\nabla\phi))$ of $\Gamma_{t}$ satisfies
$\left\|\log(p_{s}(\nabla(\phi+\xi)))-\log(p_{s}(\nabla\phi))-\frac{\langle\nabla p_{s}(\nabla\phi),\nabla\xi\rangle}{p_{s}(\nabla\phi)}\right\|_{\mathcal{C}^{0,\alpha}(\overline{\Omega})}\leq\varepsilon\|\xi\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}$ (17)
for all $\xi\in\mathbb{B}_{\delta}$. By the same means, there exists $\delta^{\prime}=\delta^{\prime}(\epsilon)>0$ such that the boundary term satisfies
$\left\|h_{s}(\nabla\phi)-\langle\nabla h_{s}(\nabla\phi),\nabla\xi\rangle\right\|_{\mathcal{C}^{0,\alpha}(\overline{\Omega})}\leq\varepsilon\|\xi\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}$ (18)
for all $\xi\in\mathbb{B}_{\delta^{\prime}}$. By choosing $\tilde{\delta}:=\min\{\delta,\delta^{\prime}\}$, (17) and (18) hold for all $\xi\in\mathbb{B}_{\tilde{\delta}}$. By the same means, there exists $\delta_{0}=\delta_{0}(\epsilon)$ such that for all $|r|<\delta_{0}$ we obtain
$\|\Gamma(s+r,\phi)-\Gamma(s,\phi)-r\,D_{\phi}\Gamma_{t}(\phi)\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq\epsilon|r|$
for all $|r|<\delta_{0}$. Since $\epsilon>0$ was arbitrary, $\Gamma$ is Fréchet differentiable over $I\times\mathcal{U}_{t}$. The continuity of the derivative can be easily deduced from Lemmas 4.1, 4.2, 4.3, and 4.4. ∎
4.2 Linearized Monge-Ampère equation and strict obliqueness
In this subsection, we start by showing that the linearized Monge-Ampère operator $D_{\phi}\Gamma_{t}(\phi_{t})$ is a second-order elliptic differential operator, see Lemma 4.8, with strictly oblique boundary conditions, see Lemma 4.9, and it is invertible as a mapping from $\mathcal{X}_{t}$ to $\mathcal{Y}_{t}$, where
$\mathcal{X}_{t}=\left\{\xi\in\mathcal{C}^{2,\alpha}(\overline{\Omega}):\int_{\partial\Omega_{t}}p_{t}\xi(\nabla\phi_{t}^{*})=0\right\},$
and
$\mathcal{Y}_{t}:=\left\{(f,g)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega):\int_{\Omega}qf=\int_{\partial\Omega_{t}}p_{t}g(\nabla\phi_{t}^{*})\right\}.$
For the sake of clarity, we recall that $p_{t}g(\nabla\phi_{t}^{*})$ is the function $x\mapsto p_{t}(x)g(\nabla\phi_{t}^{*}(x))$. Here, we equip $\mathcal{X}_{t}$ with $\mathcal{C}^{2,\alpha}$ norm and $\mathcal{Y}_{t}$ with the product norm on $\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$, under which $\mathcal{X}_{t},\mathcal{Y}_{t}$ are Banach spaces. We note that $\phi_{t}\in\mathcal{X}_{t}$, and, as shown in Theorem 4.6, $D_{t}\Gamma(\phi_{t})\xi\in\mathcal{Y}_{t}$ for every $\xi\in\mathcal{X}_{t}$. To invert it, we must obtain the existence and uniqueness of solutions in $\mathcal{X}_{t}$ to the linearized Monge-Ampère equation:
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=f\quad\rm in\,\,\Omega,$ (19)
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=g\quad\rm on\,\,\partial\Omega,$
for every $(f,g)\in\mathcal{Y}_{t}$, which is proven in Theorem 4.6. The following theorem yields the solvavilty of linearized Monge-Ampère equation and has its own value in studying the second-order elliptic partial differential equations with strictly oblique boundary conditions and a null zero-order term. To the best of our knowledge, this result is not found in the existing literature.
Theorem 4.6 (Solvability of the linearized Monge-Ampère equation).
Let the setting of Theorem 3.2 hold. Let $f\in\mathcal{C}^{0\alpha}(\overline{\Omega}),g\in\mathcal{C}^{1,\alpha}(\partial{\Omega})$. Then the linearized Monge-Ampère equation (19) admits a unique solution $\xi\in\mathcal{X}_{t}$ if and only if
$\int_{\Omega}qf=\int_{\partial\Omega_{t}}p_{t}g(\nabla\phi_{t}^{*}).$ (CC)
As a consequence, $D_{\phi}\Gamma_{t}(\phi_{t}):\mathcal{X}_{t}\to\mathcal{Y}_{t}$ is bounded and invertible.
Remark 4.7.
The condition presented in (CC) is commonly known as the compatibility condition in the literature related to the Neumann-type boundary condition.
We now proceed with the proof of Theorem 4.6. We start by proving that (19) is a second-order elliptic partial differential equation with strictly oblique boundary conditions. The following lemma controls $[D^{2}\phi_{t}]^{-1}$ from below and above yielding the uniform ellipticity of $D_{\phi}\Gamma_{t}(\phi_{t})$. The proof is omitted as it is presented in [18, Remark 1.1].
Lemma 4.8.
Let the setting of Theorem 3.2 hold, for each $t\in I$, there exists $\beta_{t}>0$ such that
$\beta_{t}^{-1}{\rm I}_{d}\leq[D^{2}\phi_{t}]^{-1}\leq\beta_{t}{\rm I}_{d}\quad\rm in\,\,\overline{\Omega}.$
The following result shows that $D_{\phi}\Gamma_{t}(\phi_{t})$ admits a strictly oblique boundary condition. Its proof follows among the lines of [39, Section 2] and it is included for the sake of completeness.
Lemma 4.9 (Strict Obliqueness).
Under the setting of Theorem 3.2, for each $t\in I$, there exist $\rho=\rho_{t}>0$ such that
$\left\langle\nabla h_{t}(\nabla\phi_{t}({x})),\nu({x})\right\rangle\geq\rho\quad\text{for all}\ x\in\partial\Omega,$
where $\nu$ denotes the unit outer normal vector-field to $\partial\Omega$.
Proof.
For brevity, we may drop any sub-indices related to $t$. Define $H(x):=h(\nabla\phi({x})),$ for ${x}\in\overline{\Omega}$ and $\chi({x}):=\left\langle H({x}),\nu({x})\right\rangle,{x}\in\partial\Omega$. Since $\phi\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$ and $h\in\mathcal{C}^{2,\alpha}(\overline{\Omega^{\prime}})$, $\chi$ is continuous on $\partial\Omega$ and admits minimum at some point ${x}_{0}\in\partial\Omega$. Let $\{{e}_{i}\}_{i=1}^{d}$ be an othonormal basis of $\mathbb{R}^{d}$. Up to a translation and rotation, we may assume ${x}_{0}={0}$, ${e}_{1},\cdots,{e}_{d-1}$ are tangential to $\partial\Omega$ at 00{0}, and $\nu({0})={e}_{d}$. Since $h(\nabla\phi)<0$ in $\Omega$ and equals zero on $\partial\Omega$, we have
$D_{\alpha}H({0})=\left\langle DH({0}),{e}_{\alpha}\right\rangle=\left\langle D^{2}\phi({0})\nabla h(\nabla\phi({0})),{e}_{\alpha}\right\rangle=D_{\alpha k}\phi D_{k}h=0\quad{\rm on}\,\,\partial\Omega,$ (20)
for $\alpha=1,\cdots,d-1,$ and
$D_{d}H({0})=\left\langle DH({0}),{e}_{d}\right\rangle=\left\langle D^{2}\phi({0})\nabla h(\nabla\phi({0})),{e}_{d}\right\rangle=D_{dk}\phi D_{k}h\geq 0\quad\rm on\,\,\partial\Omega,$ (21)
where $D_{dk}\phi:=\partial_{{e}_{d}}\partial_{{e}_{k}}\phi({0})$, $D_{k}h:=\partial_{{e}_{k}}h(D\phi({0}))$, and we are summing over repeated indices. For ease of the notations, the functions are assumed to be evaluated at 00{0} unless explicitly specified in the rest of this proof. In particular, combining (20) and (21) implies that
$D^{2}\phi\nabla h=\nabla H,$ (22)
and
$\nabla H=D_{d}H\,{e}_{d}=(0,\dots,D_{d}H)$ (23)
where $\nabla h:=(D_{1}h,\cdots,D_{d}h)^{\intercal}$. Since $\phi$ is strictly convex, $D^{2}\phi$ is invertible on $\partial\Omega$, and thus
$D_{d}h={e}_{d}^{\intercal}[D^{2}\phi]^{-1}\nabla H.$ (24)
Taking the square of both sides of (24), we obtain that
$(D_{d}h)^{2}$ $=({e}_{d}^{\intercal}[D^{2}\phi]^{-1}\nabla H)({e}_{d}^{\intercal}[D^{2}\phi]^{-1}\nabla H)$
[…]
$((D_{d}H)\in\mathbb{R})\implies$ $=(D_{d}H)^{2}{e}_{d}^{\intercal}[D^{2}\phi]^{-1}{e}_{d}{e}_{d}^{\intercal}[D^{2}\phi]^{-1}{e}_{d}$
[…]
$={e}_{d}^{\intercal}[D^{2}\phi]^{-1}{e}_{d}(\nabla h)^{\intercal}D^{2}\phi\nabla h.$
Hence,
$\chi({0})=\left\langle\nabla h(\nabla\phi({0})),\nu({0})\right\rangle=D_{d}h=\sqrt{{e}_{d}^{\intercal}[D^{2}\phi]^{-1}{e}_{d}(\nabla h)^{\intercal}D^{2}\phi\nabla h}.$
By applying Corollary 4.8, we derive the bound
$\chi({0})\geq\beta_{t}^{-2}\sqrt{{e}_{d}^{\intercal}{e}_{d}(\nabla h)^{\intercal}\nabla h},$
which implies the existence of a constant $\rho>0$ such that $\chi({0})\geq\rho$ as $\left\lVert\nabla h\right\rVert_{2}=1$ by assumption. Since $\inf_{{x}\in\partial\Omega}\chi({x})=\chi({0})$, the claim follows. ∎
Up to now, we see that the linearized Monge-Ampère operator $D_{\phi}\Gamma_{t}(\phi_{t})$ is a second-order elliptic differential operator with strictly oblique boundary condition. However, one can observe that the zero-order term in (19) is null. Then the standard Schauder theory of elliptic partial differential equations with oblique derivative boundary conditions presented in [22, Section 6.7], is not applicable. Nonetheless, as shown in the following lemma, we may prove that, if a solution exists, it is unique up to a constant shift.
Lemma 4.10.
Under the setting of Theorem 3.2, let $\xi\in\mathcal{X}_{t}$ be a solution to
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=0\quad\rm in\,\,\Omega,$
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=0\quad\rm on\,\,\partial\Omega.$
Then $\xi=0$.
Proof.
Assume that the contrary holds. By strong maximum principle [22, Theorem 3.5], $\xi$ attains its maximum at some $x_{0}\in\partial\Omega$. Let $\nu$ be the outer normal unit vector field at $\partial\Omega$. As $\Omega$ is uniformly convex, there exists a ball $B\subset\Omega$ such that $x_{0}\in\Omega$. Lemma 4.9 implies that $\left\langle\nabla h_{t}(\nabla\phi_{t}),\nu\right\rangle>0$ on $\partial\Omega$ (hence $\left\langle\nabla h_{t}(\nabla\phi_{t}(x_{0})),\nu(x_{0})\right\rangle>0$), so that Hopf’s Lemma [30, Theorem 2.5] yields that $\left\langle\nabla h_{t}(\nabla\phi_{t}(x_{0})),\nabla\xi(x_{0})\right\rangle>0,$ which contradicts the boundary condition. ∎
Now we are ready to proceed with the proof of Theorem 4.6. The uniqueness result presented in Lemma 4.10 motivates an application of the Fredholm alternative. This, combined with the approach used in [36, Theorem 3.1], establishes the necessary and sufficient conditions for the solvability of the linearized Monge-Ampère equation for a general right-hand side, specifically, for $f\in\mathcal{C}^{0,\alpha}(\overline{\Omega})$ and $g\in\mathcal{C}^{1,\alpha}(\partial{\Omega})$.
Proof of Theorem 4.6.
We prove necessity and sufficiency separately.
Necessity: Suppose $\xi\in\mathcal{X}_{t}$ solves (19), then by (5)
$\int_{\Omega}qf=\int_{\Omega}\left(p_{t}(\nabla\phi_{t})\mathrm{tr}([D^{2}\phi_{t}]^{-1}\nabla^{2}\xi)+\left\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\right\rangle\right)\det(D^{2}\phi_{t}).$ (25)
Using a change of variable from $x\mapsto\nabla\phi_{t}(x)$ and noting that $[D^{2}\phi_{t}(\nabla\phi_{t}^{*})]^{-1}=D^{2}\phi_{t}^{*}$, (25) becomes
$\int_{\Omega_{t}}\mathrm{tr}(D^{2}\phi_{t}^{*}D^{2}\xi(\nabla\phi_{t}^{*}))p_{t}+\left\langle\nabla p_{t},\nabla\xi(\nabla\phi_{t}^{*})\right\rangle=\int_{\Omega_{t}}\,\mathrm{div}(p_{t}\nabla\xi(\nabla\phi_{t}^{*})).$ (26)
By the divergence theorem, (26) can written as
$\int_{\Omega_{t}}\,\mathrm{div}(p_{t}\nabla\xi(\nabla\phi_{t}^{*}))=\int_{\partial\Omega_{t}}\left\langle p_{t}\nabla\xi(\nabla\phi_{t}^{*}),\nabla h_{t}\right\rangle=\int_{\partial\Omega_{t}}p_{t}g(\nabla\phi^{*}_{t}).$
Sufficiency: We follow the strategy presented in Theorem 3.1 in [36] to prove the result using the Fredholm alternative. By Lemma 4.8, Lemma 4.9, and [22, Theorem 6.31] the operator
$\mathbb{M}:\mathcal{C}^{2,\alpha}(\overline{\Omega})$ $\to\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$
$\xi$ $\mapsto\left(\begin{array}[]{c}L\\ B\end{array}\right)=\left(\begin{array}[]{c}{\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p(\nabla\phi_{t}),\nabla\xi\rangle}{p(\nabla\phi_{t})}\\ \langle\nabla h(\nabla\phi_{t}),\nabla\xi\rangle+\xi\end{array}\right)$
is bounded with bounded inverse $\mathbb{M}^{-1}$. The necessity proved above implies that if $\xi\in\mathcal{X}_{t}$, then
$\int_{\Omega}qL(\xi)=\int_{\partial\Omega_{t}}p_{t}[B(\xi)-\xi](\nabla\phi_{t}^{*})=\int_{\partial\Omega_{t}}p_{t}B(\xi)(\nabla\phi_{t}^{*}),$
so that $\mathbb{M}(\xi)\in\mathcal{Y}_{t}$. As a consequence, $\mathbb{M}|_{\mathcal{X}_{t}}:\mathcal{X}_{t}\to\mathcal{Y}_{t}$ is a bijection. Call $\mathcal{L}:\mathcal{Y}_{t}\to\mathcal{X}_{t}$ its (bounded) inverse and consider the equation
$\xi-\mathcal{L}(0,\xi)=\mathcal{L}(f,g).$ (27)
Here, the boundedness of $\mathcal{L}$ follows from the inverse mapping theorem [3, Corollary 2.7]. We see that $\xi\in\mathcal{X}_{t}$ solves (19) if only if $\xi$ solves (27).
Note here that $\mathcal{L}(0,\xi)$ is a well-defined element of $\mathcal{X}_{t}$ as $(0,\xi)\in\mathcal{Y}_{t}$. Call $T:\mathcal{X}_{t}\ni g\mapsto\mathcal{L}[0,g]\in\mathcal{X}_{t}$. Since $\mathcal{L}:\mathcal{Y}_{t}\to\mathcal{X}_{t}$ is bounded and
${\bf i}:\mathcal{C}^{2,\alpha}(\overline{\Omega})\cap\mathcal{X}_{t}\ni g\mapsto g\in\mathcal{C}^{1,\alpha}(\partial\Omega)\cap\left\{\xi\in\mathcal{C}^{1,\alpha}(\partial\Omega):\int_{\partial\Omega_{t}}p_{t}\xi(\nabla\phi_{t}^{*})=0\right\}$
is compact, the identity $T(g)=\mathcal{L}(0,{\bf i}(g))$ yields the compactness of $T$. Since (27) is equivalent to
$\xi-T\xi=\mathcal{L}(f,g),$ (28)
the Fredholm alternative [22, Theorem 5.3] implies that (28) admits a unique solution $\xi\in\mathcal{X}_{t}$ if and only if the homogeneous equation
$\xi-T\xi=0$ (29)
has only a trivial solution. As (29) is equivalent to (19), according to Lemma 4.10, $\xi=0$ is the unique solution of (19) in $\mathcal{X}_{t}$. The result follows. ∎
4.3 The implicit function theorem
Theorem 4.6 guarantees the invertibility of $D_{\phi}\Gamma_{t}(\phi_{t})$ between the spaces $\mathcal{X}_{t}$ and $\mathcal{Y}_{t}$. However, the range of $\Gamma$ may extend beyond $\mathcal{Y}_{t}$. Therefore, we modify the functional as follows to satisfy the compatibility conditions (CC). Set $t\in{\rm int}(I)$ and consider
$\tilde{\Gamma}^{(t)}:I\times(\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\cap\mathcal{X}_{t})$ $\to\mathcal{Y}_{t}$
$\left(\begin{array}[]{c}s\\ \phi\end{array}\right)$ $\mapsto\left(\begin{array}[]{c}\Gamma_{s}^{(1)}(\phi)-\int_{\Omega}q\Gamma_{s}^{(1)}(\phi)+\int_{\partial\Omega_{t}}p_{t}[\Gamma_{s}^{(2)}(\phi)](\nabla\phi_{t}^{*})\\ \Gamma_{s}^{(2)}(\phi)\end{array}\right).$ (34)
The following result shows that for each $t\in{\rm int}(I)$, the implicit function theorem (see e.g., [22, Theorem 17.6] or [13, Theorem 15.1 and Corollary 15.1]) can be applied over $\tilde{\Gamma}^{(t)}$.
Lemma 4.11.
Let $t\in{\rm int}(I)$ be as above. Under the assumptions of Theorem 3.2, the following holds:
  1. 1.
    $\tilde{\Gamma}^{(t)}_{s}(\phi_{s})=0$ if and only if $\Gamma_{s}(\phi_{s})=0$.
  2. 2.
    There exist an open set $I^{\prime}\times\mathcal{U}_{t}\subset I\times(\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\cap\mathcal{X}_{t})$ with $(t,\phi_{t})\in I\times\mathcal{U}_{t}$ such that $\tilde{\Gamma}^{(t)}$ is $\mathcal{C}^{1}$ in $I^{\prime}\times\mathcal{U}_{t}$.
  3. 3.
    The Fréchet derivative of $\tilde{\Gamma}^{(t)}$ w.r.t. $\phi\in\mathcal{X}_{t}$ at $(t,\phi_{t})\in I\times(\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\cap\mathcal{X}_{t})$, namely $D_{\phi}\tilde{\Gamma}_{t}(\phi_{t})$, satisfies $D_{\phi}\tilde{\Gamma}_{t}(\phi_{t})h=D_{\phi}\Gamma_{t}(\phi_{t})h$ for all $h\in\mathcal{X}_{t}$.
  4. 4.
    The bounded linear operator $D_{\phi}\tilde{\Gamma}_{t}(\phi_{t}):\mathcal{X}_{t}\to\mathcal{Y}_{t}$ is invertible.
Proof.
The second point is a direct consequence of Lemma 4.5 and the fourth point holds due to the third point and Theorem 4.6.
Proof of 1. We prove that if $\phi\in\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})$ solves
$\det(D^{2}\phi)=C\frac{q}{p_{s}(\nabla\phi)}\quad{\rm in}\,\,\Omega,\ h_{s}(\nabla\phi)=0\quad{\rm in}\,\,\partial\Omega,$
for some $C\in\mathbb{R}$, then $C=1$. First note that $C\neq 0$ as $\phi\in\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})$. In this setting it holds that
$\det(D^{2}\phi)=C\frac{q}{p_{s}(\nabla\phi)}\quad{\rm in}\,\,\Omega,\ \nabla\phi(\Omega)=\Omega_{s},$
which implies $C=\int_{\Omega}\det(D^{2}\phi)p_{s}(\nabla\phi)=\int_{\Omega_{s}}p_{s}=1$ and the claim follows.
Proof of 3. From the linearity of integration, one have for $u\in\mathcal{X}_{t}$
$D_{\phi}\tilde{\Gamma}^{(t)}_{t}(\phi_{t})(u)=\left(\begin{array}[]{c}D_{\phi}\Gamma_{t}^{(1)}(\phi_{t})(u)-\int_{\Omega}qD_{\phi}\Gamma_{t}^{(1)}(\phi_{t})(u)+\int_{\partial\Omega_{t}}p_{t}[D_{\phi}\Gamma_{t}^{(2)}(\phi_{t})(u)](\nabla\phi_{t}^{*})\\ D_{\phi}\Gamma_{t}^{(2)}(\phi_{t})(u)\end{array}\right).$
Hence the result holds after noticing that
$\int_{\Omega}qD_{\phi}\Gamma_{t}^{(1)}(\phi_{t})(u)=\int_{\partial\Omega_{t}}p_{t}[D_{\phi}\Gamma_{t}^{(2)}(\phi_{t})(u)](\nabla\phi_{t}^{*}),$
which follows from the fact that $D_{\phi}\Gamma_{t}(\phi_{t})(\mathcal{X}_{t})=\mathcal{Y}_{t}$ (see Theorem 4.6). ∎
The implicit function theorem in Banach spaces (see [13, Theorem 15.1]) implies that ${\rm int}(I)\ni t\mapsto\phi_{t}\in\mathcal{C}^{2,\alpha}$ is of class $\mathcal{C}^{1}$, meaning that $\{\phi_{t}\}_{t\in{\rm int}(I)}$ is a $\mathcal{C}^{1}$ curve. If $I$ is open, the proof finishes here. If not, Lemma 4.11 still holds for points $t$ at the boundary of $I$ by taking perturbations on the appropriate directions. Assume that $t:=\inf(I)\in I$. The mapping
$G(s,\phi)=\phi-[D_{\phi}\tilde{\Gamma}^{(t)}_{t}(\phi_{t})]^{-1}\tilde{\Gamma}^{(t)}(s,\phi)$
is a uniform contraction for values of $(s,\phi)$ close enough to $(t,\phi_{t})$. One can then conclude by the uniform contraction mapping principle that the curve $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\Omega)$ is also $\mathcal{C}^{1}$. Moreover, it holds that
$[D_{\phi}\tilde{\Gamma}_{t}^{(t)}(\phi_{t})](\partial_{t}\phi_{t})=-D_{t}\tilde{\Gamma}_{t}^{(t)}(\phi_{t}),$
so that $\partial_{t}\phi_{t}$ is the unique solution of
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=-\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}-\int_{\Omega}q\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}+\int_{\partial\Omega_{t}}p_{t}{\partial_{t}h_{t}}\quad\rm in\,\,\Omega,$ (35)
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=-\partial_{t}h_{t}(\nabla\phi_{t})\quad\rm on\,\,\partial\Omega.$
We focus now on simplifying the right hand side of (35). To do so, we prove
$\int_{\Omega_{t}}q\partial p_{t}=\int_{\partial\Omega_{t}}p_{t}{\partial_{t}h_{t}},$ (36)
which follows as a consequence of the following Reynolds transport theorem (see e.g., [17, Appendix C.4]) for implicitly defined surfaces.
Lemma 4.12.
Let $\{h_{t}\}_{t\in I}$, $\{\Omega_{t}\}_{t\in I}$ and $\Omega$ be as in Theorem 3.2. Then for every $\mathcal{C}^{1}$ curve $\{f_{t}\}_{t\in I}$ on $\mathcal{C}(\Omega^{\prime})$ it holds that
$\partial_{t}\int_{\Omega_{t}}f_{t}=\int_{\Omega^{\prime}}\partial_{t}f_{t}-\int_{\partial\Omega_{t}}f_{t}\partial_{t}h_{t}.$
Proof.
Set $t\in{\rm int}(I)$ and define the curve $T_{s}=(\nabla h_{t+s})^{-1}\nabla h_{t}$. By assumption and by the inverse function theorem, $\{T_{s}\}_{s\in I}\subset\mathcal{C}(\Omega^{\prime})$ is $\mathcal{C}^{1}$, $h_{t+s}(T_{s})=0$ in $\partial\Omega_{t}$, and $T_{s}(\Omega_{t})=\Omega_{s}$. A second-order Taylor development on $h_{t+s}$ provides the existence of a curve $\{\omega_{s}\}_{s\in I}$ on $\mathcal{C}(\partial\Omega_{t})$ such that $o(\|\omega_{s}\|_{\mathcal{C}(\partial\Omega_{t})})=o(s)$ and
$h_{t}=0=h_{t+s}(T_{s})=h_{t+s}+\langle\nabla h_{t+s},T_{s}-T_{0}\rangle+w(s)\quad\text{in}\ \partial\Omega_{t}.$
As a consequence, we get the relation
$\partial_{t}h_{t}=-\langle\left.\partial_{s}\right|_{s=0}T_{s},\nabla h_{t}\rangle\quad\text{in}\ \partial\Omega_{t}.$ (37)
Since $\nabla h_{t}$ is the unit outer-normal vector field to $\partial\Omega_{t}$, Reynolds transport theorem implies that
$\partial_{t}\int_{\Omega_{t}}f_{t}$ $=\int_{\Omega^{\prime}}\partial_{t}f_{t}+\int_{\partial\Omega_{t}}f_{t}\left\langle\left.\partial_{s}\right|_{s=0}T_{s},\nabla h_{t}\right\rangle$
and the result follows by (37). ∎