### v4

# Linearization of Monge–Ampère Equations and Statistical Applications

Abstract
Optimal transport has found numerous applications across data science, many of which require differentiating the optimal transport map with respect to the underlying probability densities in the Fréchet sense. In this work, we show that when the reference measure $Q$ is sufficiently regular in space and the curve of target measures $\{P_{t}\}_{t\in I}$ is both spatially regular and $\mathcal{C}^{1}$ in time, then the associated curve of optimal transport maps $\{\nabla\phi_{t}\}_{t\in I}$ pushing $Q$ toward $P_{t}$ is itself a $\mathcal{C}^{1}$ curve. Moreover, we identify its time derivative as the solution to the linearized Monge–Ampère equation, a second-order elliptic PDE with strictly oblique boundary conditions and a vanishing zero-order term. Our proof relies on applying the implicit function theorem to the Monge–Ampère equation with natural boundary conditions. As consequences, we establish regularity of the transport-based quantile regressor with respect to the covariates and derive a central limit theorem for smooth optimal transport maps.
1 Introduction
The optimal transport problem has become increasingly popular in data science, appearing frequently in both statistics [carlier2016vector, delBarrioQuantile2024, HallinBarrio2021, hosseini2024conditionaloptimaltransportfunction] and machine learning, including Bayesian inverse problems [hosseini2024conditionaloptimaltransportfunction, kerrigan2024dynamicconditionaloptimaltransport], generative modeling [Wgans, albergo2023stochastic], and deep learning [castin2024smoothattention], among others. In many of these applications (see, e.g., [hamm2024manifoldlearningwassersteinspace, hosseini2024conditionaloptimaltransportfunction, manole2023central, sheng2025stability, sheng2025theory]), an important and persistent challenge is to quantify how optimal transport maps change when the source measure is fixed and the target measure is perturbed. In particular, let $I\subset\mathbb{R}$ be an open interval, and let $\{P_{t}\}_{t\in I}$ denote a family of probability measures with densities $\{p_{t}\}_{t\in I}$ and supports $\{\Omega_{t}\}_{t\in I}$. The present work identifies sufficient conditions on $\{P_{t}\}_{t\in I}$ and a reference measure $Q\in\mathcal{P}(\Omega)$ guaranteeing that the curve of optimal transport maps $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\Omega)$, pushing $Q$ forward to $P_{t}$, solving
$\det(D^{2}\phi_{t})$ $=\frac{q}{p_{t}(\nabla\phi_{t})}\qquad\text{in }\Omega,$ (1)
$\nabla\phi_{t}(\Omega)$ $=\Omega_{t},$
is a $\mathcal{C}^{1}$ curve in time.
Understanding the time regularity of optimal transport maps under perturbations provides tools for establishing central limit theorems [del2025distributional, manole2023central] and for analyzing the statistical properties of plug-in estimators [Manole2024Plugin, balakrishnan2025stability, balakrishnan2025statistical]. Moreover, such regularity can be used to study the smoothness of velocity fields in manifold learning problems in the Wasserstein space [hamm2024manifoldlearningwassersteinspace].
Although the spatial regularity of $(t,x)\mapsto\phi(t,x):=\phi_{t}(x)$ is well understood from Caffarelli’s regularity theory (see [Caffarelli1990, caffarelli1992boundary, delBarrioAdvNon2024, figalli2017monge, Figalli2018]), the time regularity is subtler. In [loeper2005regularity, Proposition 4.1], the author shows that if the supports $\{\Omega_{t}\}_{t\in I}$ are constant in $t$ and both the densities and supports are $\mathcal{C}^{\infty}$, then $\phi$ is differentiable in $t$. As in the present work, the proof relies on an application of the implicit function theorem. Furthermore, in [loeper2006fully], the same author uses this result to establish the existence and uniqueness of classical solutions to the semigeostrophic equations on a three-dimensional flat torus.
More recent results in [manole2023central] also focus on the setting where both source and target measures are supported on the flat torus $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$, and the transport maps are periodic (see [CORDEROERAUSQUIN1999199]). Their analysis relies crucially on the periodic boundary conditions of the Monge–Ampère equation on the torus, and therefore does not extend to the setting of interest here, particularly when the supports $\{\Omega_{t}\}_{t\in I}$ vary in time.
A related but distinct line of work concerns stability of optimal transport maps, with extensive literature on both qualitative and quantitative aspects. For example, qualitative results include: Theorem 4.2 in [philippis2013regularity], which states that if $p_{t^{\prime}}\to p_{t}$ in $L^{1}(\Omega^{\prime})$, then $\nabla\phi_{t^{\prime}}\to\nabla\phi_{t}$ in $W^{1,\gamma}_{\mathrm{loc}}(\Omega)$ for some $\gamma>1$; Corollary 5.23 in [villani2009optimal], which shows that weak convergence $p_{t^{\prime}}\to p_{t}$ implies $\nabla\phi_{t^{\prime}}\to\nabla\phi_{t}$ in measure; and [Segers.2022], which proves uniform convergence $\|\nabla\phi_{t^{\prime}}-\nabla\phi_{t}\|_{L^{\infty}}\to 0$ under additional regularity assumptions. Quantitative stability estimates can be found in [manole2021plugin, delalande2021quantitative]. However, stability results establish only continuity of the optimal transport map under perturbations and do not address finer differentiability properties. In contrast, identifying the time derivative of the transport map is central to many of the aforementioned applications.
Overall, there are few results concerning the time regularity of optimal transport maps when the target measures $\{P_{t}\}_{t\in I}$ have time-varying supports $\{\Omega_{t}\}_{t\in I}$.
Results and Methodology
In this work, we show that if the reference measure $Q$ is sufficiently regular in space, and the curve of probability measures $\{P_{t}\}_{t\in I}$ is sufficiently regular in space and evolves smoothly in time (in the sense of Definition 3.1), then the corresponding curve of optimal transport maps $\{\nabla\phi_{t}\}_{t\in I}$ also evolves smoothly in time. Moreover, its time derivative solves the linearized Monge–Ampère equation, which is a second-order elliptic partial differential equation with strictly oblique boundary conditions and a null zero-order term (see Theorem 3.2).
Our approach follows the strategy of linearizing the Monge–Ampère equation (see [figalli2017monge, Villani2003]) and applying the implicit function theorem. To handle the time-varying supports, we make essential use of the tool introduced in [urbas1997second], which characterizes the evolution of the supports of the target measures via convex defining functions.
The proof of our main result relies on analyzing the following functional derived from the Monge–Ampère equation:
$\Gamma(t,\phi)=\begin{pmatrix}\Gamma_{t}^{(1)}(\phi)\\[4.0pt] \Gamma_{t}^{(2)}(\phi)\end{pmatrix}=\begin{pmatrix}\log(\det(D^{2}\phi))-\log(q)+\log(p_{t}(\nabla\phi))\\[4.0pt] h_{t}(\nabla\phi)\end{pmatrix}$,
where $h_{t}$ is a uniformly convex defining function for ${\rm supp}(P_{t})$. Analyzing the invertibility of its Fréchet derivative leads to the study of the following second-order elliptic PDE:
${\rm tr}\!\big([D^{2}\phi_{t}]^{-1}D^{2}\xi\big)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=f\quad\text{in }\Omega,$
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=g\quad\text{on }\partial\Omega.$
In Theorem 5.6, we show that this PDE is solvable between the spaces
$\mathcal{X}_{t}=\left\{\xi\in\mathcal{C}^{2,\alpha}(\overline{\Omega}):\int_{\partial\Omega_{t}}p_{t}\,\xi(\nabla\phi_{t}^{*})=0\right\}$
and
$\mathcal{Y}_{t}=\left\{(f,g)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega):\int_{\Omega}qf=\int_{\partial\Omega_{t}}p_{t}\,g(\nabla\phi_{t}^{*})\right\}.$
The proof then proceeds by applying the implicit function theorem to a re-centered version of $\Gamma$, since the range of $\Gamma$ generally exceeds the space $\mathcal{Y}_{t}$.
Building on our theoretical results, we present two statistical applications. The first concerns the transport-based regression problem (Section 4.1), where the conditional distribution $P_{Y\mid X=x}$ plays the role of the time-varying target measure. We show that the associated transport-based quantile regressor varies smoothly with respect to the covariate $x$. Our second application (Section 4.2) establishes a central limit theorem for the smoothed optimal transport map, whose limiting distribution is again characterized through the linearized Monge–Ampère equation.
Organization
The remainder of the paper is organized as follows: Section 2 introduces and recalls key notations used throughout for the reader’s convenience. In Section 3, we state the main result and its assumptions. Section 4 applies our theoretical results to various statistical problems. Section 5 is devoted to the proof of the main result.
2 Notation
For any Borel probability measure $P$ over $\mathbb{R}^{d}$ (i.e., $P\in\mathcal{P}(\mathbb{R}^{d})$), let $\mathrm{supp}(P)$ denote its topological support, which is defined to be $\mathrm{supp}(P):=\mathbb{R}^{d}\setminus\bigcup\{U\subseteq\mathbb{R}^{d}:U\text{ is open },P(U)=0\}.$ The support of a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is defined to be $\mathrm{supp}(f):=\overline{\{x\in\mathbb{R}^{d}:f(x)\neq 0\}}$, where the overline the notes the topological closure. The interior of a set $A$ in a topological space $B$ is denoted as ${\rm int}(A)$. We call a measure absolutely continuous if it is absolutely continuous with respect to the Lebesgue measure $\ell_{d}$. The same applies with densities, unless the contrary is stated, a density of a probability measure $P$ is the Radon–Nikodym derivative of $P$ w.r.t. $\ell_{d}$. The integration, of a function $f$ is always w.r.t. $\ell_{d}$ (or the corresponding Hausdorff measure $\mathcal{H}^{k}$ if a $k$-dimensional surface $S$ were involved) and we simply write $\int f=\int f{\rm d}\ell_{d}=\int f(x){\rm d}x$ (or $\int_{S}f=\int_{S}f{\rm d}\mathcal{H}^{k}$ for the surface case). The indicator function of a Borel measurable set $A$ is denoted as ${\bf 1}_{A}$.
Let $U\subseteq\mathbb{R}^{d}$ be bounded, $k\in\mathbb{N},\alpha\in[0,1]$, we denote by $\mathcal{C}^{k,\alpha}(U)$ the space consisting of functions whose $k$-th order partial derivatives are uniformly Hölder continuous with exponent $\alpha$. We adopt the convention that $\mathcal{C}^{k,0}(U)=\mathcal{C}^{k}(U)$. Let $f\in\mathcal{C}^{1}(U)$, $e\in\mathbb{R}^{d}$, we denote by $\partial_{e}u$ the directional derivative along the direction $e$ and denote by $\nabla f:=(\partial_{e_{1}}f,\dots,\partial_{e_{d}}f)^{\intercal}$, where $\{e_{j}\}_{j=1}^{d}$ denotes the canonical basis of $\mathbb{R}^{d}$. Hessian matrix of a function $f\in\mathcal{C}^{2}(U)$ is denoted as $D^{2}f$. Let $F:\mathbb{R}^{d}\to\mathbb{R}^{d}$, we denote by $\mathrm{div}(F)$ the divergence of $F$. Denote by $\mathbb{R}_{+}^{d}$ the half-space $\{x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}:x_{d}>0\}$. We call domain every open, bounded, connected, and non-empty subset $\Omega$ of $\mathbb{R}^{d}$. Moreover, $\Omega$ is said to be a $\mathcal{C}^{k,\alpha}$ domain $(k\in\mathbb{N},\alpha\in[0,1])$ if for every $p\in\partial\Omega$, the boundary of $\Omega$, there exists a neighborhood $B=B(x_{0})$ of $x_{0}$ in $\mathbb{R}^{d}$ and a diffeomorphism $\Psi:B\rightarrow D:=\Psi(B)\subset\mathbb{R}^{d}$ such that
$\text{(i) $\Psi(B\cap\Omega)\subset\mathbb{R}_{+}^{d}$; \quad(ii) $\Psi(B\cap\partial\Omega)\subset\partial\mathbb{R}_{+}^{d}$; \quad(iii) $\Psi\in\mathcal{C}^{k,\alpha}(B),\Psi^{-1}\in\mathcal{C}^{k,\alpha}(D)$}.$
As curve $\{b_{t}\}_{t\in I}$ in a Banach space $(B,\|\cdot\|_{B})$ is $\mathcal{C}^{1}$ if there exists a continuous function $\partial_{t}b_{t}:I\to B$ such that the limit
$\lim_{h\to 0}\left\|\frac{1}{h}(b_{t+h}-b_{t})-\partial_{t}b_{t}\right\|_{B}=0,\quad\text{for all}\,\,t\in I\,.$ (2)
A mapping $F:\mathcal{U}\to\mathcal{Y}$, were $(\mathcal{X},\|\cdot\|_{\mathcal{X}})$ and $(\mathcal{Y},\|\cdot\|_{\mathcal{Y}})$ are Banach spaces and $\mathcal{U}\subset\mathcal{X}$ is open, is said to be Fréchet differentiable at $x\in\mathcal{U}$ if there exists a bounded linear functional $A(x):\mathcal{X}\to\mathcal{Y}$ such that
$\lim_{h\to 0}\frac{\left\lVert F(x+h)-F(x)-A(x)h\right\rVert_{\mathcal{Y}}}{\left\lVert h\right\rVert_{\mathcal{X}}}=0$
for every $h\in\mathcal{X}$. We say that $F$ is Fréchet differentiable in $\mathcal{U}$ if $F$ is Fréchet differentiable at every point $x\in\mathcal{U}$. We say that $F$ is $\mathcal{C}^{1}$ in $\mathcal{U}$ if it is Fréchet differentiable in $\mathcal{U}$ and the mapping $A:\mathcal{X}\to\mathcal{L}(\mathcal{X},\mathcal{Y})$, where $\mathcal{L}(\mathcal{X},\mathcal{Y})$ denotes the space of bounded linear functional from $\mathcal{X}$ to $\mathcal{Y}$, is continuous w.r.t. the norm topology. We denote by ${\rm tr}(M)$ the trace of a matrix $M$. Given two symmetric matrices $A,B$ with dimension $d\times d$ over $\mathbb{R}$, we say that $A\leq B$ if BAB-A is positive semi-definite. For quantities $a,b$, we write $a\lesssim b$ if there exists a constant $C>0$ (which may depend on other parameters depending on context) such that $a\leq Cb$.
3 Main Result
In this section, we state the main result and introduce its assumptions. We consider a curve $\{P_{t}\}_{t\in I}$ of probability measures with both smooth (and uniformly convex) supports and densities.
As mentioned in the introduction, our main result accommodates time-dependent supports. Due to this dependency, the curve $\{p_{t}\}_{t\in I}$ is not well-defined in $\mathcal{C}^{1,\gamma}(\Omega^{\prime})$ as in the case $P_{t}={\rm Unif}[0,1+t]$ for $t\in(0,1)$. To address this issue, we will start assume that $\{p_{t}\}_{t\in I}$ has a $\mathcal{C}^{1,\gamma}$ extension. Later, in Lemma 3.4, we will provide sufficient conditions for this extension.
To begin with, we define our notion of a $\mathcal{C}^{1}$ curve of probability measures in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ for a given set $\Omega^{\prime}$. The definition below captures two aspects of the curve of probability measures: changes in support and changes in the density functions.
Definition 3.1.
A curve $\{P_{t}\}_{t\in I}$ of probability measures over a domain $\Omega^{\prime}$ is said to be $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ if the following conditions are satisfied:
  1. (1)
    There exists $\kappa\in(0,+\infty)$ and a $\mathcal{C}^{1}$ curve of convex functions $\{h_{t}\}_{t\in I}$ in $\mathcal{C}^{2,\alpha}(\overline{\Omega^{\prime}})$ such that
    $\Omega_{t}:={\rm int}(\mathrm{supp}(P_{t}))=\left\{y\in\mathbb{R}^{d}:h_{t}(y)<0\right\},$
    with $\|\nabla h_{t}\|=1$ on $\partial\Omega_{t}$, $\bigcup_{t\in I}\Omega_{t}+\frac{1}{\kappa}\mathbb{B}:=\left\{y:{\rm dist}\left(y,\bigcup_{t\in I}\Omega_{t}\right)<\frac{1}{\kappa}\right\}\subset\Omega^{\prime}$, and
    $\frac{1}{\kappa}{\rm I}_{d}\leq D^{2}h_{t}\leq\kappa{\rm I}_{d}\quad\text{in }\Omega^{\prime}.$
  2. (2)
    There exists a $\mathcal{C}^{1}$ curve $\{\log(p_{t})\}_{t\in I}$ in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ such that $P_{t}=p_{t}{\bf 1}_{\Omega_{t}}$.
The function $h_{t}$ is commonly known as the convex defining function of $\mathrm{supp}(P_{t})$. It is noteworthy that $\Omega_{t}=\left\{y\in\mathbb{R}^{d}:h_{P}(y)<0\right\}$ represents an open, bounded, and uniformly $\mathcal{C}^{2,\alpha}$ convex domain. Conversely, given a uniformly convex set $\Omega_{t}$, one can construct the function $h_{t}$ to be smooth and uniformly convex, approximating $\text{dist}(\cdot,\partial\Omega_{t})+\frac{1}{2}\text{dist}(\cdot,\partial\Omega_{t})^{2}$ near $\partial\Omega_{t}$. For example, if $\Omega_{t}=\mathbb{B}(0,1)=\{x\in\mathbb{R}^{d}:\|x\|=1\}$, one might choose $h_{P}(y):=\frac{1}{2}(\|y\|^{2}-1)$ (see, e.g., [figalli2017monge, pp 40]).
Theorem 3.2.
Let $\Omega$ be a $\mathcal{C}^{2}$ uniformly convex domain and let $\Omega^{\prime}\subset\mathbb{R}^{d}$ be a domain. If the curve $\{P_{t}\}_{t\in I}$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ with $0<\gamma\leq 1$ (in the sense of Definition 3.1) and let $\log(q)\in\mathcal{C}^{\alpha}(\overline{\Omega})$ with $0<\alpha<\gamma$, then the curve $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\overline{\Omega})$, where $\phi_{t}$ solves (1), is also $\mathcal{C}^{1}$. Moreover, the time derivative of $I\ni t\mapsto\phi_{t}\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$, denoted as $\partial_{t}\phi_{t}$, is the unique solution (up to an additive shift) of the linearized Monge–Ampère equation
[…]
Remark 3.3.
  1. (i)
    As we will see in Lemma 5.2, requiring the curve $\{P_{t}\}_{t\in I}$ to be $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ (rather than merely in $\mathcal{C}^{1,\alpha}$) is necessary to ensure the continuity of the map $\phi\mapsto\nabla p_{t}\circ\phi$. This level of regularity is also assumed in [urbas1997second].
  2. (ii)
    Recent advances on the Monge–Ampère equation [chen2021global] suggest that the assumptions on $\Omega$ and $\{\Omega_{t}\}_{t\in I}$ could be relaxed to convexity and $\mathcal{C}^{1,1}$ boundary regularity. However, to avoid additional technical complications in our arguments, we adopt the classical assumptions used in [caffarelli1996boundary, urbas1997second], which are sufficient to cover the primary applications we have in Section 4.
  3. (iii)
    The positivity of $\alpha$ is crucial for the application of Schauder theory in [gilbarg1983trudinger], and therefore cannot be relaxed using our current techniques.
In Theorem 3.2, we implicitly assume that each density $p_{t}$, originally defined only on $\Omega_{t}$, admits an extension to the fixed domain $\Omega^{\prime}$. The following result provides a sufficient condition ensuring the existence of such an extension, thereby justifying the applicability of Theorem 3.2. We present its proof below, while the proof of Theorem 3.2 is deferred to Section 5.
Lemma 3.4.
Assume that $I$ is open. Let $\{h_{t},\Omega_{t}\}_{t\in I}$ and $\Omega^{\prime}$ be defined as in Definition 3.1, let $h:(t,x)\mapsto h(t,x)=h_{t}(x)$ be $\mathcal{C}^{1,\alpha}$ over $I\times\Omega^{\prime}$, and let $\log(p):(t,x)\mapsto\log(p(t,x))=\log(p_{t}(x))$ be a $\mathcal{C}^{1,\alpha}$ function over the set $\mathcal{D}:=\{(t,x):\ t\in I,\ x\in\overline{\Omega}_{t}\}.$ Then the curve $\{P_{t}\}_{t\in I}$ with $P_{t}=p_{t}{\bf 1}_{\Omega_{t}}$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\alpha}(\overline{\Omega^{\prime}})$ in the sense of Definition 3.1.
Proof.
We want to extend $\log p$ from ${\rm int}(\mathcal{D}):={\rm int}\left(\{(t,x):\,t\in I,\ x\in\Omega_{t}\}\right)$ to $I\times\Omega^{\prime}$ using [gilbarg1983trudinger, Lemma 6.37]. Since $I$ is open and connected, we may assume without loss of generality that $I=\mathbb{R}$. Then, for every $(t_{0},x_{0})\in\partial\mathcal{D}$, it suffices to find a neighborhood $B(t_{0},x_{0})$ of $(t_{0},x_{0})$ and a $\mathcal{C}^{1,\alpha}$ diffeomorphism $\Psi:B(t_{0},x_{0})\to D:=\Psi(B(t_{0},x_{0}))\subset\mathbb{R}^{d+1}$ such that $\Psi(B(t_{0},x_{0})\cap\mathcal{D})\subset\mathbb{R}^{d+1}_{+}$, $\Psi(B(t_{0},x_{0})\cap\partial\mathcal{D})\subset\partial\mathbb{R}^{d+1}_{+}$, and $\Psi\in\mathcal{C}^{1,\alpha}(B(t_{0},x_{0})),\Psi^{-1}\in\mathcal{C}^{1,\alpha}(D)$. A candidate map is
$\Psi:\mathbb{R}^{d+1}\ni(t,x_{1},\dots,x_{d-1},x_{d})\mapsto(t,x_{1},\dots,x_{d-1},-h_{t}(x))\in\mathbb{R}^{d+1}.$
Since $h_{t}$ is a convex defining function of $\Omega_{t}$, the first two requirements are easily satisfied, and $\Psi$ is $\mathcal{C}^{1,\alpha}$. We now show that $\Psi$ is one-to-one in a small neighbourhood of $(t_{0},x_{0})$. As $\{x\in\Omega:h_{t}(x)<0\}$ is open and $h_{t}$ is uniformly convex, one may choose $\delta$ small enough such that $\|\nabla h_{t}(x)\|\neq 0$ for all $(t,x)\in B:=\mathbb{B}((t_{0},x_{0}),\delta)$. We proceed with the proof by contradiction. Suppose there exists $(t,x)=(t,x_{1},\dots,x_{d})\neq(s,y)=(s,y_{1},\dots,y_{d})\in B$ such that $\Psi(t,x)=\Psi(s,y)$, then
$t=s,\,\,x_{i}=y_{i},\quad i=1,\dots,d-1,$
and $h_{t}(x)=h_{t}(y)$. Define $\tilde{h}:z\mapsto h_{t}(x_{z})$, where $x_{z}:=(x_{1},\dots,x_{d-1},z)$, then $\tilde{h}$ is also $\mathcal{C}^{1,\alpha}$ uniformly convex. Moreover, as $|\tilde{h}^{\prime}(z)|\neq 0$ over $\mathcal{Z}:=\{z:(t,x_{z})\in B\}$, $\tilde{h}^{\prime}(z)=\left\langle\nabla h_{t}(x_{z}),e_{d}\right\rangle$ is either positive or negative over $\mathcal{Z}$ by the monotonicity of the gradient of a uniformly convex function. Therefore, the fundamental theorem of calculus yields that
$0=\tilde{h}(y_{d})-\tilde{h}(x_{d})=(y_{d}-x_{d})\int_{0}^{1}\tilde{h}^{\prime}(x_{d}+t(y_{d}-x_{d}))\,dt\neq 0,$
which leads to a contradiction. ∎
Under the assumption that the supports $\{\Omega_{t}\}_{t\in I}$ remain constant in $t$, we no longer require that $\{P_{t}\}_{t\in I}$ forms a $\mathcal{C}^{1}$ curve in the sense of Definition 3.1; it suffices to be a $\mathcal{C}^{1}$ curve in the standard sense. In this case, we recover the classical result of [loeper2005regularity]. It is worth noting that the following theorem is already sufficient to derive asymptotic confidence intervals for the optimal transport map, as in [manole2023central].
Corollary 3.5.
Let $\Omega$ (resp. $\Omega^{\prime}$) be a $\mathcal{C}^{2}$ (resp. $\mathcal{C}^{2,\gamma}$ with $0<\gamma\leq 1$) uniformly convex domain. Let $\log(p_{t})\in\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ and let $\log(q)\in\mathcal{C}^{\alpha}(\overline{\Omega})$ with $0<\alpha<\gamma$ such that $1=\int_{\Omega}q=\int_{\Omega^{\prime}}p_{t}=1$ for all $t\in I$. If the curve $\{\log(p_{t})\}_{t\in I}\subset\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ is $\mathcal{C}^{1}$ (in the sense of (2)), then the curve $\{\nabla\phi_{t}\}_{t\in I}\subset\mathcal{C}^{1,\alpha}(\overline{\Omega})$, where $\phi_{t}$ solves (1) for $\Omega_{t}=\Omega^{\prime}$, is also $\mathcal{C}^{1}$. Moreover, the time derivative of $I\ni t\mapsto\phi_{t}\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$, denoted as $\partial_{t}\phi_{t}$, is the unique solution (up to an additive shift) of the linearized Monge–Ampère equation
[…]
where $\nu_{\Omega^{\prime}}$ denotes the unit outer normal vector-field to $\partial\Omega^{\prime}$.
Corollary 3.5 follows immediately from Theorem 3.2 once we show that the family $\{\log(p_{t})\}_{t\in I}\subset\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ can be extended to a larger set $\Omega^{\prime}_{\kappa}:=\Omega^{\prime}+\frac{1}{\kappa}\mathbb{B}$ for some $\kappa$ depending on the uniform convexity of $\Omega^{\prime}$ (see Lemma 3.6 below). This extension is crucial for computing the Fréchet derivative of the target functional (see Lemma 5.5). In contrast with Lemma 3.4, the fact that the support does not depend on $t$ eliminates the main challenge to our analysis. As a consequence, we no longer need to impose any extra regularity on the curve $t\mapsto\log(p_{t})$.
Lemma 3.6.
Under the assumptions of Corollary 3.5, the curve $\{P_{t}\}_{t\in I}$, with $P_{t}=p_{t}{\bf 1}_{\Omega^{\prime}}$, is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ (in the sense of Definition 3.1).
Proof.
Fix $t\in I$. Since $\Omega^{\prime}$ is $\mathcal{C}^{2,\gamma}$, for each $x\in\partial\Omega^{\prime}$, there exists a ball $B(x)=\mathbb{B}(x,\epsilon_{x})$ and a $\mathcal{C}^{2,\gamma}$ diffeomorphism $\Psi=\Psi_{x}:B(x)\to D_{x}:=\Psi_{x}(B(x))\subset\mathbb{R}^{d}$ that straightens the boundary (see (2)). Since $\partial\Omega^{\prime}\subset\bigcup_{x\in\partial\Omega^{\prime}}\mathbb{B}\left(x,\frac{\epsilon_{x}}{2}\right)$ and $\partial\Omega^{\prime}$ is compact, there exists a finite covering $\partial\Omega^{\prime}\subset\bigcup_{i=1}^{n}\mathbb{B}\left(x_{i},\frac{\epsilon_{x_{i}}}{2}\right)$, with $x_{i}\in\partial\Omega^{\prime}$. Let $\kappa>0$ be such that $\Omega^{\prime}\subset\Omega_{\kappa}^{\prime}\subset\bigcup_{i=1}^{n}\mathbb{B}\left(x_{i},\frac{\epsilon_{x_{i}}}{2}\right)\cup\Omega^{\prime}$.
Denote by $u_{t}:=\log p_{t}$ for ease of notation. Set $i\in\{1,\dots,n\}$ and
$D_{+}^{(i)}:=\Psi(\Omega^{\prime}\cap B(x_{i}))=D_{x_{i}}\cap\mathbb{R}_{+}^{d}.$
Setting $\tilde{u}_{t}^{(i)}(y):=u\circ\Psi^{-1}_{x_{i}}(y)$ for every $y=(y_{1},\dots,y_{d})=:(y^{\prime},y_{d})\in D_{+}^{(i)}$, the extension of $\tilde{u}_{t}$ to $D_{x_{i}}$ is given by
$\tilde{w}_{t}^{(i)}(y^{\prime},y_{d}):=-3\tilde{u}_{t}^{(i)}(y^{\prime},-y_{d})+4\tilde{u}_{t}^{(i)}(y^{\prime},-y_{d}/2),\quad y_{d}<0,$
so that $\tilde{w}_{t}^{(i)}\in\mathcal{C}^{1,\gamma}(D_{x_{i}})$ (see [gilbarg1983trudinger, Lemma 6.37]) and
[…]
is a $\mathcal{C}^{1,\gamma}$ extension of $u_{t}$ in $B(x_{i})\cup\Omega^{\prime}$. Let $\{\eta_{i}\}_{i=1}^{n}$ be a locally finite partition of unity subordinate to the covering $\{B(x_{i})\}_{i=1}^{n}$. Then we have
$w_{t}=\sum_{i=1}^{n}\eta_{i}w_{t}^{(i)}\in\mathcal{C}^{1,\gamma}(\overline{\Omega_{\kappa}^{\prime}})$
and $w_{t}=u_{t}$ in $\overline{\Omega^{\prime}}$ for all $t\in I$. Since $u_{t}=\log(p_{t})$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}})$ in the standard sense, Lemma 5.2 (below) shows that for each $i$, the curve $\{w_{t}^{(i)}\}_{t\in I}$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}\left(\overline{\mathbb{B}\left(x_{i},\epsilon_{x_{i}}/2\right)\cup\Omega^{\prime}}\right)$, which a fortiori implies that $\{w_{t}\}_{t\in I}$ is $\mathcal{C}^{1}$ in $\mathcal{C}^{1,\gamma}(\overline{\Omega^{\prime}_{\kappa}})$. This concludes the proof. ∎
4 Applications
4.1 Transport-based Quantile Regression
Our first application concerns the transport-based quantile regression problem [carlier2016vector, delBarrioQuantile2024]. Let $(X,Y)\in\mathbb{R}^{m+d}$ be two random vectors with joint probability distribution $P_{X,Y}$. The transport-based quantile regressor $(u,x)\mapsto\nabla_{u}\phi(u,x)=\mathbb{Q}_{Y\mid X}(u,x)$ of random vectors $Y$ on $X$ is defined as the unique Borel function such that, for $P_{X}$-almost every $x$, the function $u\mapsto\nabla_{u}\phi(u,x)$ is the transport-based quantile function of $P_{Y|X=x}$—the conditional probability measure of $Y$ given $X=x$ —which we assume to have density $p_{Y|X}(\cdot,x)$ [Chernozhukov2017, HallinBarrio2021, ghosal2022multivariate, Shi2022]. Recall that transport-based quantile function of a probability measure $P$ as the unique gradient of a convex function pushing a fixed reference measure $Q$ (with density $q$ w.r.t. the Lebesgue measure) forward to $P$. Therefore, for each $x\in{\rm int}({\rm supp}(P_{X}))$, $u\mapsto\phi(u,x)$ solves the conditional Monge–Ampère equation [carlier2016vector]:
$\det(D^{2}_{u}\phi(\cdot,x))$ $=\frac{q}{p_{Y|X}(\nabla_{u}\phi(\cdot,x),x)}\quad{\rm in}\ \Omega,$
$\nabla\phi(\Omega\times\{x\})$ $={\rm supp}(P_{Y|X}(\cdot,x)).$
In this setting, our goal translates into studying the regularity of the transport-based quantile regressor $(u,x)\mapsto\nabla_{u}\varphi(u,x)$ with respect to the covariates $x$.
Figure 1: Multivariate quantile regression (two-dimensional variable of interest $\mathbf{Y}=(Y_{1},Y_{2})^{\intercal}$; univariate regressor $X$), showing the conditional medians (red) and the conditional quantile contours of order $\tau=0.2$ (black), $\tau=0.4$ (green), $\tau=0.8$ (yellow). The number of samples generated is $n=320,050$. The nonparametric estimator is that of [delBarrioQuantile2024], trivially adapted to the change of reference measure.
Figure 1 illustrates the expressive capabilities of the transport-based quantile regressor in visualizing complex data. It effectively captures nonlinear trends, heteroskedasticity, and the overall shape of the conditional distributions. In the case $m=1$ and $d=2$, we visualize the regression median (red) together with quantile tubes of orders $\tau=0.2$ (black), $0.4$ (green), and $0.8$ (yellow), defined via the conditional quantile contours, i.e., the image of the mapping $x\mapsto\bigl(x,\mathbb{Q}_{Y\mid X}\bigl(\partial\mathbb{B}(0,\sqrt{\tau}),x\bigr)\bigr)$, for the model
$\mathbf{Y}=\begin{pmatrix}Y_{1}\\ Y_{2}\end{pmatrix}=R(X)\begin{pmatrix}\frac{1}{2}(X^{2}+1)Z_{1}\\ -\frac{1}{4}X^{2}+\frac{1}{2}(X^{2}+1)Z_{2}\end{pmatrix}$,
where $X\sim{\rm Unif}[-2,2]$ and
$R(X)=\begin{pmatrix}\cos\left(\frac{\pi}{2}(2-X)\right)&-\sin\left(\frac{\pi}{2}(2-X)\right)\\ \sin\left(\frac{\pi}{2}(2-X)\right)&\cos\left(\frac{\pi}{2}(2-X)\right)\end{pmatrix}\in\mathbb{R}^{2\times 2},\qquad\mathbf{Z}:=(Z_{1},Z_{2})^{\top}=\tilde{\mathbf{Z}}\,{\bf 1}_{\{\|\tilde{\mathbf{Z}}\|\leq 100\}}.$
The random vector $\tilde{\mathbf{Z}}$ follows a four-component Gaussian mixture, where each component has equal weight, common covariance matrix $\Sigma=10\,{\rm I}_{d}$, and means $^{\top}, (0.866,0.5)(-0.866,-0.5)^{\top}$, $(0,0)^{\top}$, and $(0,1)^{\top}$, respectively. The reference measure for the quantile regressor in Figure 1 is the uniform distribution on the unit ball.
From Theorem 3.2 and Lemma 3.4, we obtain as a corollary the desired regularity of the transport-based quantile regressor, which in particular apply to the toy example above. This smoothness can in turn be used to develop asymptotic theory [KoenkerBook] and to perform causal and counterfactual inference (see, e.g., [ImbensNewey2009, Chernozhukov2013]), as in the classical one-dimensional case.
Corollary 4.1.
Let $\Omega$, $\Omega^{\prime}$ and the reference measure $Q={\bf 1}_{\Omega}qdx$ be as in Theorem 3.2. Let $(X,Y)\in\mathbb{R}^{m+d}$ be a pair of random variables with probability law $P_{X,Y}=P_{Y|X}P_{X}$. Assume that
  1. (1)
    for each $x\in{\rm int}({\rm supp}(P_{X}))$, the conditional probability measure has support within $\Omega^{\prime}$ and the log density $(x,y)\mapsto\log p_{Y|X}(y,x)$ is $\mathcal{C}^{1,\gamma}$ over $\overline{{\rm supp}(P_{X,Y})}$ and
  2. (2)
    the convex defining function $h(\cdot,x)$ of ${\rm supp}(P_{Y\mid X}(\cdot,x)))$ is $\mathcal{C}^{1,\gamma}$ over $\Omega^{\prime}\times{\rm supp}(P_{X})$.
Then the function ${\rm int}({\rm supp}(P_{x}))\ni x\mapsto\mathbb{Q}_{Y|X}(\cdot,x)\in\mathcal{C}^{1,\alpha}(\overline{\Omega})$ is $\mathcal{C}^{1}$. In particular, $\nabla_{x}\mathbb{Q}_{Y|X}\in\mathcal{C}(\overline{\Omega}\times{\rm int}({\rm supp}(P_{X})))$.
4.2 Central Limit Theorem for Smooth Optimal Transport Map
The smooth optimal transport problem is introduced as a way to alleviate the curse of dimensionality in estimating the Wasserstein distance [Goldfeld2020, goldfeld2020gaussian, goldfeld2020asymptotic, nietert2021smooth, GoldfeldKato2024]. Let $\Omega$ be a $\mathcal{C}^{2,1}$ uniformly convex domain. Let $K:\overline{\Omega}\times\overline{\Omega}\to(0,\infty)$ be a $\mathcal{C}^{2,1}$ kernel satisfying $\int_{\Omega}K(x,y)\,{\rm d}y=1$ and $\int_{\Omega}K(x,y)\,{\rm d}x=1$ for all $(x,y)\in\Omega\times\Omega$. Let the reference measure $Q\in\mathcal{P}(\Omega)$ satisfy $\log(q)\in\mathcal{C}^{\alpha}(\overline{\Omega})$ for some $0<\alpha<1$. The smooth (or convoluted) optimal transport problem between $P\in\mathcal{P}(\Omega)$ and $Q$ is defined as
$\min_{\pi\in\Pi(P_{K},Q)}\int\|x-y\|^{2}\,d\pi(x,y),$ (4)
where $P_{K}$ admits the smoothed density $\int K(\cdot,x)\,dP(x)$.
Suppose we observe i.i.d. samples $X_{1},\dots,X_{n}$ from $P$ and define the empirical measure $\widehat{P}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. Let $T=\nabla\phi$ be the optimal transport map from $Q$ to $P_{K}$, and let $\widehat{T}_{n}$ denote the optimal transport map from $Q$ to $\widehat{P}_{n,K}$. In this section, we establish a central limit theorem for the smooth optimal transport maps $\widehat{T}_{n}$ (see Proposition 4.4).
As expected, the result follows from the functional delta method. To apply it, one needs the Hadamard differentiability of the mapping $\Phi$ that sends a density $p$ to the optimal transport map $\nabla\phi_{q\to p}$ pushing $q$ forward to $p$. This differentiability follows from our general result Corollary 3.5, together with the lemma below. We remark that this lemma provides a general principle showing that smoothness along every $\mathcal{C}^{1}$ curve implies Hadamard differentiability, which may be of independent interest to the reader.
Lemma 4.2.
Let $(\mathcal{X},\|\cdot\|_{\mathcal{X}})$, $(\mathcal{Y},\|\cdot\|_{\mathcal{Y}})$ be normed spaces. Assume that $\mathcal{K}\subset\mathcal{X}$ is open. Let $\Phi:\mathcal{K}\to\mathcal{Y}$ be a functional. Suppose that for any $\mathcal{C}^{1}$ curve $\eta:[0,1]\mapsto\mathcal{K}$ the function $\Phi\circ\eta:[0,1]\to\mathcal{Y}$ is also $\mathcal{C}^{1}$. Then $\Phi$ is Hadamard differentiable at any $x_{0}\in\mathcal{K}$. That is, for any $t_{k}\to 0^{+}$ and $h_{k}\to h\in\mathcal{X}$ with $\{h_{k}\}_{k\geq 0}\subset\mathcal{X}$, it follows that
$\lim_{k\to\infty}\frac{\Phi(x_{0}+t_{k}h_{k})-\Phi(x_{0})}{t_{k}}=D\Phi(x_{0})(h)=\frac{d}{dt}\bigg|_{t=0}\Phi(x_{0}+ht).$
Proof.
For any $h\in\mathcal{X}$, the curve $x_{t}=x_{0}+th$ is $\mathcal{C}^{1}$, for $t$ small enough, then there exists a mapping $D\Phi(x_{0}):\mathcal{X}\to\mathcal{Y}$ such that
$\lim_{t\to 0^{+}}\left\|\frac{\Phi(x_{t})-\Phi(x_{0})}{t}-D\Phi(x_{0})(h)\right\|_{\mathcal{Y}}=0\,.$
That is, $\Phi$ is Gateaux differentiable at $x_{0}$. Note that $\Phi$ is positive homogeneous as for $\lambda>0$,
$D\Phi(x)(\lambda h)=\lim_{t\to 0^{+}}\frac{\Phi(x_{0}+t\lambda h)-\Phi(x_{0})}{t}=\lim_{s\to 0^{+}}\frac{\Phi(x_{0}+sh)-\Phi(x_{0})}{s/\lambda}=\lambda D\Phi(x)(h)\,.$
We proceed by contradiction. Suppose that
$\lim_{k}\left\|\frac{\Phi(x_{0}+t_{k}h_{k})-\Phi(x_{0})}{t_{k}}-D\Phi(x_{0})(h)\right\|_{\mathcal{Y}}=\alpha\in(0,\infty]\,.$
for some sequence $t_{k}\to 0^{+}$ and $h_{k}\to h\in\mathcal{X}$ with the convention that $\alpha=\infty$ if the sequence $\left\{\|(\Phi(x_{0}+t_{k}h_{k})-\Phi(x_{0}))/t_{k}\|_{\mathcal{Y}}\right\}_{k\geq 1}$ diverges. The above condition shows that
$\lim_{k}\left\|\frac{\Phi(x_{0}+t_{k}h_{k})-\Phi(x_{0}+t_{k}h)}{t_{k}}\right\|_{\mathcal{Y}}>0\,.$ (5)
Define $\eta_{k}:s\mapsto x_{0}+t_{k}(h+s(h_{k}-h))\in\mathcal{K}$ for $k$ large enough, then $\Phi\circ\eta_{k}$ is $\mathcal{C}^{1}$ in $\mathcal{Y}$. Hence, the mean-value theorem (see e.g., [Lang.functional.93, Theorem 4.2.]) and the positive homogeneity of $D\Phi$ give that
$\left\|\frac{\Phi(x_{0}+t_{k}h_{k})-\Phi(x_{0}+t_{k}h)}{t_{k}}\right\|_{\mathcal{Y}}$ $\leq\sup_{s\in[0,1]}\|D\Phi(\eta_{k}(s))\left(h_{k}-h\right)\|_{\mathcal{Y}}$ (6)
$=\|h_{k}-h\|_{\mathcal{X}}\sup_{s\in[0,1]}\left\|D\Phi(\eta_{k}(s))\left(\frac{h_{k}-h}{\|h_{k}-h\|_{\mathcal{X}}}\right)\right\|_{\mathcal{Y}}\,.$
Write $v_{k}=\frac{h_{k}-h}{\|h_{k}-h\|_{\mathcal{X}}}$. If $\{\|D\Phi(\eta_{k}(s))(v_{k})\|_{\mathcal{Y}}:k\in\mathbb{N},\,s\in[0,1]\}$ is bounded, then (6) yields that
$\left\|\frac{\Phi(x_{0}+t_{k}h_{k})-\Phi(x_{0}+t_{k}h)}{t_{k}}\right\|_{\mathcal{X}}\lesssim\|h_{k}-h\|_{\mathcal{X}}\to 0,$
contradicting (5). Therefore, as $h_{k}\to h$ and $t_{k}\to 0^{+}$, after taking subsequences, we can choose a sequence $\{x_{n}\}_{n\geq 1}\subset\mathcal{X}$ such that $\|x_{n}-x_{0}\|_{\mathcal{X}}\leq 2^{-n}$ and $\|D\Phi(x_{n})(v_{n})\|_{\mathcal{Y}}\geq 2^{n^{2}}$. Applying [Kriegl1997, Corollary 2.10] shows that there exist a $\mathcal{C}^{\infty}$ curve $\eta:[-1,1]\to\mathcal{X}$ and two sequences of positive numbers $\{s_{n}\}_{n\geq 1}\subset(0,\infty)$ and $\{s_{n}^{\prime}\}_{n\geq 1}\subset(0,\infty)$ with $s_{n},s_{n}^{\prime}\to 0$ and such that $\eta(s_{n}+t)=x_{n}+t\frac{v_{n}}{2^{n}}$ for $t\in(-s_{n}^{\prime},s_{n}^{\prime})$. Since $\Phi\circ\eta$ is $\mathcal{C}^{1}$ by assumption, we derive that
$\frac{d}{dt}\bigg|_{t=0}\Phi(\eta(s_{n}+t))=\frac{1}{2^{n}}D\Phi(x_{n})(v_{n}).$ (7)
Moreover, $\left\|\frac{d}{dt}\bigg|_{t=0}\Phi(\eta(s_{n}+t))\right\|_{\mathcal{Y}}$ is finite for all $n\in\mathbb{N}$. On the other hand, (7) and the fact that $\|D\Phi(x_{n})(v_{n})\|_{\mathcal{Y}}\geq 2^{n^{2}}$ imply
$\left\|\frac{d}{dt}\bigg|_{t=0}\Phi(\eta(s_{n}+t))\right\|_{\mathcal{Y}}=\frac{1}{2^{n}}\left\|D\Phi(x_{n})(v_{n})\right\|_{\mathcal{Y}}>2^{n^{2}-n}\to\infty,$
which leads to a contradiction. ∎
As a consequence, we arrive at the following result.
Corollary 4.3.
The functional $\Phi:\mathcal{C}^{2}(\overline{\Omega})\to\mathcal{C}^{1,\alpha}(\overline{\Omega})$, mapping a density $p\in\mathcal{C}^{2}(\overline{\Omega})$ to its transport map $\nabla\phi_{q\to p}$ from $q$ to $p$, is Hadamard differentiable at $P_{K}$ tangentially to
$\mathcal{X}=\left\{\text{$f\in\mathcal{C}^{2}(\overline{\Omega})$ with $\int_{\Omega}f=0$}\right\}.$
Proof.
Since $P_{K}$ is uniformly bounded away from zero, there exists an open neighborhood $\mathcal{U}\subset\mathcal{X}$ of zero such that $P+f$ is a probability density. The function $\mathcal{U}\ni f\mapsto\Phi(P_{K}+f)$ is differentiable along curves in the sense of Lemma 4.2. Hence, applying Corollary 3.5 concludes the result. ∎
Now, we ready to show the central limit theorem. Recall that for any Banach space $(\mathcal{X},\|\cdot\|_{\mathcal{X}})$, we say that a sequence of $\mathcal{X}$-valued random elements converges weakly to $X$, write as $X_{n}\overset{d}{\to}X$, if for any bounded continuous function $f:\mathcal{X}\to\mathbb{R}$, $\mathbb{E}[f(X_{n})]\to\mathbb{E}[f(X)]$.
Proposition 4.4.
It follows that
$\sqrt{n}\left(\widehat{T}_{n}-T\right)\xrightarrow{d}\nabla\tilde{\mathbb{G}}$ (8)
in $\mathcal{C}^{1,\alpha}(\overline{\Omega})$, where $\tilde{\mathbb{G}}$ solves the following equation a.e.,
${\rm tr}([D^{2}\phi]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{K}(T),\nabla\xi\rangle}{p_{K}(T)}$ $=-\frac{\tilde{\mathbb{G}}}{p_{K}(T)}\quad\rm in\,\,\Omega\,$,
$\langle\nabla h_{t}(T),\nabla\xi\rangle$ $=0\quad\rm on\,\,\partial\Omega\,$
and $\tilde{\mathbb{G}}$ is the centered tight Gaussian process in $\mathcal{C}^{2}(\overline{\Omega})$ with covariance function
$\mathbb{E}[\tilde{\mathbb{G}}(x)\tilde{\mathbb{G}}(y)]={\rm Cov}\left(K(y,X_{1})K(x,X_{1})\right)\,.$
Proof.
In view of $\widehat{P}_{n,K}=\frac{1}{n}\sum_{i=1}^{n}K(\cdot,X_{i})\in\mathcal{C}^{2,1}(\overline{\Omega})$ and as $\mathcal{C}^{2,1}(\overline{\Omega})$ norm of $K(\cdot,X_{1})$ is bounded by a deterministic constant $C$, the following CLT in $\mathcal{C}^{2}(\overline{\Omega})$ (see e.g., [Naresh.CLT.1976, Theorem 3.2]) holds
$\sqrt{n}\left(\widehat{P}_{n,K}-{P}_{K}\right)\xrightarrow{d}\mathbb{G}.$ (9)
Then the results follows from (9), Corollary 4.3 and the functional delta-method [VaartWellner2023, Theorem 3.10.4]. ∎
5 Proof of Theorem 3.2
In this section, we provide the proof of Theorem 3.2. We recall that for each $t$, the function $\phi_{t}:\Omega\to\mathbb{R}$ is strictly convex and solves the boundary problem (1). Following [urbas1997second], we rewrite (1) in a form that allows for a linearization:
$\det(D^{2}\phi_{t})$ $=\frac{q}{p_{t}(\nabla\phi_{t})}\quad{\rm in}\,\,\Omega,$ (10)
$h_{t}(\nabla\phi_{t})$ $=0\quad{\rm in}\,\,\partial\Omega.$
As indicated previously, the strategy is to differentiate (10) and apply the implicit function theorem [deimling2013nonlinear, Theorem 15.1] over the functional
$\Gamma:I\times\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})$ $\to\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$ (11)
$\left(\begin{array}[]{c}t\\ \phi\end{array}\right)$ $\mapsto\left(\begin{array}[]{c}\Gamma_{t}^{(1)}(\phi)\\ \Gamma_{t}^{(2)}(\phi)\end{array}\right)=\left(\begin{array}[]{c}\log(\det(D^{2}\phi))-\log(q)+\log(p_{t}(\nabla\phi))\\ h_{t}(\nabla\phi)\end{array}\right),$
where for $\lambda\geq 0$,
$\mathcal{C}^{k,\alpha}_{>\lambda}(\bar{\Omega}):=\{f\in\mathcal{C}^{k,\alpha}(\overline{\Omega}):D^{2}f>\lambda I\}.$
We observe that the terms $\log(p_{t}(\nabla\phi))$ and $h_{t}(\nabla\phi)$ appear in the definition of $\Gamma$. Therefore, we must ensure that the composition of Hölder continuous functions with the same exponents remains Hölder continuous with exponent unchanged. Unfortunately, such a claim is untrue in general. For example, if $f,g\in\mathcal{C}^{0,\alpha}(\overline{\Omega})$, then $f\circ g\in\mathcal{C}^{0,\alpha^{2}}(\overline{\Omega})$ rather than being $\mathcal{C}^{0,\alpha}(\overline{\Omega})$. However, the claim is valid when either $f$ or $g$ has higher regularity as in the following lemma, which is an immediate result from [de1999regularity, Theorem 4.3].
Lemma 5.1 (Compositions of Hölder Functions).
For any $\alpha,\beta\in[0,1]$ and integers $l\geq k\geq 0$ with $l\geq 1$, there exists a constant $C=C(k,l,\alpha,\beta,\Omega,\Omega^{\prime})>0$ such that for all $f\in\mathcal{C}^{k,\alpha}(\Omega^{\prime})$ and $g\in\mathcal{C}^{l+1,\beta}(\overline{\Omega})$ satisfying $\nabla g(\overline{\Omega})\subset\Omega^{\prime}$, we have $f\circ\nabla g\in\mathcal{C}^{k,\min(\alpha,\beta)}(\overline{\Omega})$ and
$\|f\circ\nabla g\|_{\mathcal{C}^{k,\min(\alpha,\beta)}(\overline{\Omega})}\leq C\bigl(1+\|g\|_{\mathcal{C}^{l+1,\beta}(\overline{\Omega})}^{k+\alpha}\bigr)\|f\|_{\mathcal{C}^{k,\alpha}(\Omega^{\prime})}.$
As a direct consequence of Lemma 5.1, we establish that $\Gamma$ is well-defined. We conclude this subsection by presenting a set of properties related to Hölder spaces, which will be instrumental in the subsequent subsection for proving that the Fréchet derivative of $\Gamma$ with respect to $\phi$ is well-defined. The following lemma, adapted from [de1999regularity, Proposition 6.1,6.2], states the stability of Hölder functions under composition, and justifies the condition in Theorem 3.2, where we assume that $p_{t}$ is $\mathcal{C}^{1,\gamma}$ rather than merely $\mathcal{C}^{1,\alpha}$.
Lemma 5.2.
Under the same setting as in Lemma 5.1, the map
$\mathcal{C}^{k,\alpha}(\Omega^{\prime})\ni f\mapsto f\circ\nabla g\in\mathcal{C}^{k,\min{(\alpha,\beta)}}(\overline{\Omega})$
is linear and continuous. In addition, if $\alpha>\beta$, then the map
$\mathcal{C}^{l+1,\beta}(\overline{\Omega})\ni g\mapsto f\circ\nabla g\in\mathcal{C}^{k,\beta}(\overline{\Omega})$
is also continuous. More precisely, there exists $\delta,\rho,M>0$ such that if $\left\lVert g^{\prime}-g\right\rVert_{\mathcal{C}^{l+1,\beta}(\overline{\Omega})}<\delta$, one has that
$\|f\circ\nabla g-f\circ\nabla g^{\prime}\|_{\mathcal{C}^{k,\beta}(\overline{\Omega})}\leq M\|f\|_{\mathcal{C}^{k,\alpha}(\Omega^{\prime})}\|g-g^{\prime}\|_{\mathcal{C}^{l+1,\beta}(\overline{\Omega})}^{\rho}.$
The next result is standard and can be found in [gilbarg1983trudinger, Section 4.1].
Lemma 5.3 (Product of Hölder Functions).
For any $\alpha,\beta\in[0,1],l\geq k\geq 0$, there exists a constant $C=C(k,l,\alpha,\beta,{\rm diam}(\Omega))>0$ such that for all $f\in\mathcal{C}^{k,\alpha}(\overline{\Omega}),g\in\mathcal{C}^{l,\beta}(\overline{\Omega})$,
$\left\lVert fg\right\rVert_{\mathcal{C}^{l,\min(\alpha,\beta)}(\overline{\Omega})}\leq C\left\lVert f\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}\left\lVert g\right\rVert_{\mathcal{C}^{l,\beta}(\overline{\Omega})}.$
As a consequence of Lemma 5.1 and Lemma 5.3, we immediately arrive at the following result as the function $1/g$ may be viewed as the composition of $x\mapsto 1/x$ and $g$, and $x\mapsto 1/x$ is $\mathcal{C}^{\infty}$ on $\lambda,\infty)$.
Lemma 5.4.
Let $f,g\in\mathcal{C}^{k,\alpha}(\overline{\Omega})$ and $g\geq c_{0}$ for some constant $c_{0}>0$, then $f/g\in\mathcal{C}^{k,\alpha}(\overline{\Omega})$ and there exists $C=C(k,\alpha,c_{0},{\rm diam}(\Omega))$ such that
$\left\lVert f/g\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}\leq C\left\lVert f\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}(1+\left\lVert g\right\rVert_{\mathcal{C}^{k,\alpha}(\overline{\Omega})}^{k+\alpha}).$
5.1 Fréchet Derivative of the Functional
Now we show that the map $\Gamma$ is $\mathcal{C}^{1}$. The proof follows the standard approach used in showing the openness in the continuity method for the Monge–Ampère equation with Dirichlet boundary conditions (see [figalli2017monge]). Due to the time-varying supports $\Omega_{t}$, there are technical challenges in ensuring that the functional is always well-defined within its respective spaces, as addressed in the second statement in the following lemma.
Lemma 5.5.
Under the assumptions of Theorem 3.2, the following holds:
  1. (1)
    $\Gamma_{t}(\phi_{t})=0$ for all $t\in I$.
  2. (2)
    For each $t\in I$ there exists an open set $I^{\prime}\times\mathcal{U}_{t}\subset I\times\mathcal{C}^{2,\alpha}_{>0}(\overline{\Omega})$ with $(t,\phi_{t})\in I^{\prime}\times\mathcal{U}_{t}$ such that $\Gamma$ is $\mathcal{C}^{1}$ in $I^{\prime}\times\mathcal{U}_{t}$.
  3. (3)
    For every $t\in I$, the Fréchet derivative of $\Gamma$ at $(t,\phi_{t})$ w.r.t. $\phi$ is the linear functional
    $D_{\phi}\Gamma_{t}(\phi_{t}):\mathcal{C}^{2,\alpha}(\overline{\Omega})$ $\to\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$
    $\xi$ $\mapsto\left(\begin{array}[]{c}{\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}\\ \langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle\end{array}\right),$
    which we call the linearized Monge–Ampère operator at $(t,\phi_{t})$.
  4. (4)
    For every $t\in I$, the Fréchet derivative of $\Gamma$ at $(t,\phi_{t})$ w.r.t. $t$ is the element
    $D_{t}\Gamma_{t}(\phi_{t})=\left(\begin{array}[]{c}\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}\\ \partial_{t}h_{t}(\nabla\phi_{t})\end{array}\right)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)\,.$
Proof.
The first claim is straightforward by (1), $\nabla\phi_{t}(\partial\Omega)=\partial\Omega_{t}$, and $h_{t}(\partial\Omega_{t})=\{0\}$. We now prove the rest of the claims together. Fix $t\in I$. The boundary regularity provided in [caffarelli1996boundary, urbas1997second] yields the existence of a constant $C=C_{t}$ such that
$\|\phi_{t}\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq C\quad{\rm and}\quad\|\phi_{t}^{*}\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq C,$
which implies that $D^{2}\phi_{t}$ belongs to the set of (strictly) positive definite $d\times d$ real matrices over the field $\mathbb{R}$ (see also Lemma 5.8 below). Denote such a set as $\mathcal{M}^{+}_{d\times d}$. Choose $\delta_{t}>0$ such that $D^{2}(\phi_{t}+\xi)\in\mathcal{M}^{+}_{d\times d}$ and $\nabla(\phi_{t}+\xi)(\overline{\Omega})\subset\Omega^{\prime}$ for all
$\xi\in\mathbb{B}_{\delta_{t}}:=\left\{f\in\mathcal{C}^{2,\alpha}(\overline{\Omega}):\|f\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq\delta_{t}\right\}.$
(Note that the latter is possible due to the assumption $\bigcup_{t\in I}\Omega_{t}+\frac{1}{\kappa}\mathbb{B}\subset\Omega^{\prime}$.) Set $\mathcal{U}_{t}:=(\phi_{t}+\mathbb{B}_{\delta_{t}})\subseteq\mathcal{C}^{2,\alpha}_{>0}(\overline{\Omega})$. We show the second claim with the set $I\times\mathcal{U}_{t}$. Fix $(s,\phi)\in I\times\mathcal{U}_{t}$. It is well-known (see, e.g., [figalli2017monge, Section 3]) that
$\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\supset\mathcal{U}_{t}\ni\phi\mapsto\log\det(D^{2}\phi)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})$
is $\mathcal{C}^{1}$ with directional derivative
$\frac{d}{d\varepsilon}\Big|_{\varepsilon=0}\log\det(A+\varepsilon B)={\rm tr}(A^{-1}B).$
Moreover, as
$\Gamma(s,\phi)=\left(\begin{array}[]{c}-\log(q)+\log(p_{s}(\nabla\phi))\\ h_{s}(\nabla\phi)\end{array}\right)+\left(\begin{array}[]{c}\log(\det(D^{2}\phi))\\ 0\end{array}\right)$
and $\log(q)$ does not vary with $(s,\phi)$, it suffices to prove the claims for
$(s,\phi)\mapsto(\log(p_{s}(\nabla\phi)),h_{s}(\nabla\phi)).$
Fix $\epsilon>0$. The chain rule [Shapiro1990Directional, Proposition 3.6] and Lemma 5.1 imply the existence $\delta=\delta(\epsilon)>0$ such that the term $\phi\mapsto\log(p_{t}(\nabla\phi))$ of $\Gamma_{t}$ satisfies
$\left\|\log(p_{s}(\nabla(\phi+\xi)))-\log(p_{s}(\nabla\phi))-\frac{\langle\nabla p_{s}(\nabla\phi),\nabla\xi\rangle}{p_{s}(\nabla\phi)}\right\|_{\mathcal{C}^{0,\alpha}(\overline{\Omega})}\leq\varepsilon\|\xi\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}$ (12)
for all $\xi\in\mathbb{B}_{\delta}$. By the same means, there exists $\delta^{\prime}=\delta^{\prime}(\epsilon)>0$ such that the boundary term satisfies
$\left\|h_{s}(\nabla(\phi+\xi))-h_{s}(\nabla\phi)-\langle\nabla h_{s}(\nabla\phi),\nabla\xi\rangle\right\|_{\mathcal{C}^{1,\alpha}(\partial{\Omega})}\leq\varepsilon\|\xi\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}$ (13)
for all $\xi\in\mathbb{B}_{\delta^{\prime}}$. Therefore, by choosing $\tilde{\delta}:=\min\{\delta,\delta^{\prime}\}$, we derive that for all $\xi\in\mathbb{B}_{\tilde{\delta}}$,
$\left\|\Gamma(s,\phi+\xi)-\Gamma(s,\phi)-D_{\phi}\Gamma_{s}(\phi)(\xi)\right\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial{\Omega})}\leq\varepsilon\|\xi\|_{\mathcal{C}^{0,\alpha}(\overline{\Omega})}\,.$
Similarly, there exists $\delta_{0}=\delta_{0}(\epsilon)$ such that for all $|r|<\delta_{0}$, the following holds
$\left\|\Gamma(s+r,\phi)-\Gamma(s,\phi)-r\,D_{t}\Gamma_{s}(\phi)\right\|_{\mathcal{C}^{2,\alpha}(\overline{\Omega})}\leq\epsilon|r|$
for all $|r|<\delta_{0}$. Since $\epsilon>0$ was arbitrary, we conclude that $\Gamma$ is Fréchet differentiable over $I\times\mathcal{U}_{t}$. Finally, the continuity of the derivatives $(s,\phi)\mapsto(D_{\phi}\Gamma_{s}(\phi),D_{t}\Gamma_{s}(\phi))$ can be easily deduced from Lemmas 5.1, 5.2, 5.3, and 5.4. ∎
5.2 Linearized Monge–Ampère Equation
In this subsection, we start by showing that the linearized Monge–Ampère operator $D_{\phi}\Gamma_{t}(\phi_{t})$ is a second-order elliptic differential operator (see Lemma 5.8) with strictly oblique boundary conditions (see Lemma 5.9). As a consequence, it is invertible (see Theorem 5.6) as a mapping from $\mathcal{X}_{t}$ to $\mathcal{Y}_{t}$, where
$\mathcal{X}_{t}=\left\{\xi\in\mathcal{C}^{2,\alpha}(\overline{\Omega}):\int_{\partial\Omega_{t}}p_{t}\xi(\nabla\phi_{t}^{*})=0\right\},$
and
$\mathcal{Y}_{t}:=\left\{(f,g)\in\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega):\int_{\Omega}qf=\int_{\partial\Omega_{t}}p_{t}g(\nabla\phi_{t}^{*})\right\}.$
Here, we recall that $p_{t}g(\nabla\phi_{t}^{*})$ is the function $x\mapsto p_{t}(x)g(\nabla\phi_{t}^{*}(x))$. Here, we equip $\mathcal{X}_{t}$ with $\mathcal{C}^{2,\alpha}$ norm and $\mathcal{Y}_{t}$ with the product norm on $\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$, under which $\mathcal{X}_{t},\mathcal{Y}_{t}$ are Banach spaces. We note that $\phi_{t}\in\mathcal{X}_{t}$, and, as shown in Theorem 5.6, $D_{t}\Gamma(\phi_{t})\xi\in\mathcal{Y}_{t}$ for every $\xi\in\mathcal{X}_{t}$. To invert it, we must obtain the existence and uniqueness of solutions in $\mathcal{X}_{t}$ to the linearized Monge–Ampère equation:
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=f\quad\rm in\,\,\Omega,$ (14)
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=g\quad\rm on\,\,\partial\Omega,$
for every $(f,g)\in\mathcal{Y}_{t}$.Due to the absence of a zero-order term, the standard Schauder theory for elliptic equations with oblique derivative boundary conditions [gilbarg1983trudinger, Section 6.7] does not apply directly. Motivated by the Fredholm-type argument used in [nardi2015schauder, Theorem 3.1] to establish solvability of the Neumann problem, and by the strategy of [delanoe1991classical] for solving the second boundary-value problem for the Monge–Ampère equation in the presence of a non-vanishing zero-order term when $d=2$, we obtain the following theorem. It establishes solvability of the linearized Monge–Ampère equation and is of independent interest for second-order elliptic equations with strictly oblique boundary conditions and a vanishing zero-order term.
Theorem 5.6 (Solvability of the Linearized Monge–Ampère Equation).
Under the setting of Theorem 3.2. Let $f\in\mathcal{C}^{0,\alpha}(\overline{\Omega}),g\in\mathcal{C}^{1,\alpha}(\partial{\Omega})$. Then the linearized Monge–Ampère equation (14) admits a unique solution $\xi\in\mathcal{X}_{t}$ if and only if
$\int_{\Omega}qf=\int_{\partial\Omega_{t}}p_{t}g(\nabla\phi_{t}^{*}).$ ($C$)
As a consequence, $D_{\phi}\Gamma_{t}(\phi_{t}):\mathcal{X}_{t}\to\mathcal{Y}_{t}$ is bounded and invertible.
Remark 5.7.
The condition presented in (CC) is commonly known as the compatibility condition in the literature related to the Neumann-type boundary condition.
We now proceed with the proof of Theorem 5.6. We start by proving that (14) is a second-order elliptic partial differential equation with strictly oblique boundary conditions. The following lemma controls $D^{2}\phi_{t}]^{-1}$ from below and above yielding the uniform ellipticity of $D_{\phi}\Gamma_{t}(\phi_{t})$. The proof is omitted as it is presented in [figalli2017monge, Remark 1.1].
Lemma 5.8.
Let the setting of Theorem 3.2 hold, for each $t\in I$, there exists $\beta_{t}>0$ such that
$\beta_{t}^{-1}{\rm I}_{d}\leq[D^{2}\phi_{t}]^{-1}\leq\beta_{t}{\rm I}_{d}\quad\rm in\,\,\overline{\Omega}.$
The following result shows that $D_{\phi}\Gamma_{t}(\phi_{t})$ admits a strictly oblique boundary condition. Its proof follows among the lines of [urbas1997second, Section 2].
Lemma 5.9 (Strict Obliqueness).
Under the setting of Theorem 3.2, for each $t\in I$, there exist $\rho_{t}>0$ such that
$\left\langle\nabla h_{t}(\nabla\phi_{t}(x)),\nu(x)\right\rangle\geq\rho_{t}\quad\text{for all}\ x\in\partial\Omega,$
where $\nu$ denotes the unit outer normal vector-field to $\partial\Omega$.
Proof.
For brevity, we may drop any sub-indices related to $t$. Define $H(x):=\nabla h(\nabla\phi(x))$ for $x\in\overline{\Omega}$ and $\chi(x):=\left\langle H(x),\nu(x)\right\rangle$ for $x\in\partial\Omega$. Since $\phi\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$ and $h\in\mathcal{C}^{2,\alpha}(\overline{\Omega^{\prime}})$, $\chi$ is continuous on $\partial\Omega$ and admits minimum at some point $x_{0}\in\partial\Omega$. Let $\{{e}_{i}\}_{i=1}^{d}$ be an orthonormal basis of $\mathbb{R}^{d}$. Up to a translation and rotation, we may assume $x_{0}={0}$, ${e}_{1},\dots,{e}_{d-1}$ are tangential to $\partial\Omega$ at 0{0}, and $\nu({0})=e_{d}$. Since $h(\nabla\phi)<0$ in $\Omega$ and equals zero on $\partial\Omega$, we have
$D_{j}H({0})=\left\langle DH({0}),e_{j}\right\rangle=\left\langle D^{2}\phi({0})\nabla h(\nabla\phi({0})),e_{j}\right\rangle=D_{jk}\phi D_{k}h=0\quad{\rm on}\,\,\partial\Omega,$ (15)
for $j=1,\dots,d-1,$ and
$D_{d}H({0})=\left\langle DH({0}),e_{d}\right\rangle=\left\langle D^{2}\phi({0})\nabla h(\nabla\phi({0})),e_{d}\right\rangle=D_{dk}\phi D_{k}h\geq 0\quad\rm on\,\,\partial\Omega,$ (16)
where $D_{dk}\phi:=\partial_{e_{d}}\partial_{{e}_{k}}\phi({0})$, $D_{k}h:=\partial_{{e}_{k}}h(D\phi({0}))$, and we are summing over repeated indices. For ease of the notations, the functions are assumed to be evaluated at 0{0} unless explicitly specified in the rest of this proof. In particular, combining (15) and (16) implies that
$D^{2}\phi\nabla h=DH,$ (17)
and
$DH=D_{d}H\,e_{d}=(0,\dots,D_{d}H)$ (18)
where $\nabla h=(D_{1}h,\dots,D_{d}h)^{\intercal}$. Since $\phi$ is strictly convex, $D^{2}\phi$ is invertible on $\partial\Omega$, and thus
$D_{d}h=e_{d}^{\intercal}[D^{2}\phi]^{-1}DH.$ (19)
Taking the square of both sides of (19), we obtain that
$(D_{d}h)^{2}$ $=(e_{d}^{\intercal}[D^{2}\phi]^{-1}DH)(e_{d}^{\intercal}[D^{2}\phi]^{-1}DH)$
[…]
$=e_{d}^{\intercal}[D^{2}\phi]^{-1}e_{d}(e_{d}D_{d}H)^{\intercal}[D^{2}\phi]^{-1}(D_{d}He_{d})$
[…]
$=e_{d}^{\intercal}[D^{2}\phi]^{-1}e_{d}(\nabla h)^{\intercal}D^{2}\phi\nabla h.$
Hence,
$\chi({0})=\left\langle\nabla h(\nabla\phi({0})),\nu({0})\right\rangle=D_{d}h=\sqrt{e_{d}^{\intercal}[D^{2}\phi]^{-1}e_{d}(\nabla h)^{\intercal}D^{2}\phi\nabla h}.$
By applying Lemma 5.8, we derive the bound
$\chi({0})\geq\beta_{t}^{-2}\sqrt{e_{d}^{\intercal}e_{d}\|\nabla h\|^{2}},$
which implies the existence of a constant $\rho_{t}>0$ such that $\chi({0})\geq\rho_{t}$ as $\|\nabla h\|=1$ by assumption. Since $\inf_{x\in\partial\Omega}\chi(x)=\chi({0})$, the claim follows. ∎
So far, we have shown that the linearized Monge–Ampère operator $D_{\phi}\Gamma_{t}(\phi_{t})$ is a second-order elliptic differential operator with strictly oblique boundary condition. Moreover, as shown in the following lemma, we may prove that, if a solution exists, it is unique up to a constant shift.
Lemma 5.10.
Under the setting of Theorem 3.2, let $\xi\in\mathcal{X}_{t}$ be a solution to
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=0\quad\rm in\,\,\Omega,$
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=0\quad\rm on\,\,\partial\Omega.$
Then $\xi=0$.
Proof.
Assume that the contrary holds. By strong maximum principle [gilbarg1983trudinger, Theorem 3.5], $\xi$ attains its maximum at some $x_{0}\in\partial\Omega$. Let $\nu$ be the outer normal unit vector field at $\partial\Omega$. As $\Omega$ is uniformly convex, there exists a ball $B\subset\Omega$ such that $x_{0}\in\Omega$. Lemma 5.9 implies that $\left\langle\nabla h_{t}(\nabla\phi_{t}),\nu\right\rangle>0$ on $\partial\Omega$ (hence $\left\langle\nabla h_{t}(\nabla\phi_{t}(x_{0})),\nu(x_{0})\right\rangle>0$), so that Hopf’s Lemma [han2011elliptic, Theorem 2.5] yields that $\left\langle\nabla h_{t}(\nabla\phi_{t}(x_{0})),\nabla\xi(x_{0})\right\rangle>0,$ which contradicts the boundary condition. ∎
Now we are ready to proceed with the proof of Theorem 5.6. As mentioned earlier, the uniqueness result presented in Lemma 5.10 motivates an application of the Fredholm alternative. It remains to identify the necessary conditions for the solvability of the linearized Monge–Ampère equation for a general $f\in\mathcal{C}^{0,\alpha}(\overline{\Omega})$ and $g\in\mathcal{C}^{1,\alpha}(\partial{\Omega})$.
Proof of Theorem 5.6.
We prove necessity and sufficiency separately.
Necessity: Suppose $\xi\in\mathcal{X}_{t}$ solves (14), then by (1)
$\int_{\Omega}qf=\int_{\Omega}\left(p_{t}(\nabla\phi_{t})\mathrm{tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\left\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\right\rangle\right)\det(D^{2}\phi_{t})\,.$ (20)
Using a change of variable from $x\mapsto\nabla\phi_{t}(x)$ and noting that $D^{2}\phi_{t}(\nabla\phi_{t}^{*})]^{-1}=D^{2}\phi_{t}^{*}$, (20) becomes
$\int_{\Omega_{t}}\mathrm{tr}(D^{2}\phi_{t}^{*}D^{2}\xi(\nabla\phi_{t}^{*}))p_{t}+\left\langle\nabla p_{t},\nabla\xi(\nabla\phi_{t}^{*})\right\rangle=\int_{\Omega_{t}}\,\mathrm{div}(p_{t}\nabla\xi(\nabla\phi_{t}^{*})).$ (21)
By the divergence theorem [evans2022partial, Appendix C.1], (21) can be written as
$\int_{\Omega_{t}}\,\mathrm{div}(p_{t}\nabla\xi(\nabla\phi_{t}^{*}))=\int_{\partial\Omega_{t}}\left\langle p_{t}\nabla\xi(\nabla\phi_{t}^{*}),\nabla h_{t}\right\rangle=\int_{\partial\Omega_{t}}p_{t}g(\nabla\phi^{*}_{t}).$
Sufficiency: By Lemma 5.8, Lemma 5.9, and [gilbarg1983trudinger, Theorem 6.31], the operator
$\mathbb{M}:\mathcal{C}^{2,\alpha}(\overline{\Omega})$ $\to\mathcal{C}^{0,\alpha}(\overline{\Omega})\times\mathcal{C}^{1,\alpha}(\partial\Omega)$
$\xi$ $\mapsto\left(\begin{array}[]{c}L\\ B\end{array}\right)=\left(\begin{array}[]{c}{\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}\\ \langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle+\xi\end{array}\right)$
is bounded with bounded inverse $\mathbb{M}^{-1}$. The necessary condition above implies that if $\xi\in\mathcal{X}_{t}$, then
$\int_{\Omega}qL(\xi)=\int_{\partial\Omega_{t}}p_{t}[B(\xi)-\xi](\nabla\phi_{t}^{*})=\int_{\partial\Omega_{t}}p_{t}B(\xi)(\nabla\phi_{t}^{*}),$
so that $\mathbb{M}(\xi)\in\mathcal{Y}_{t}$. As a consequence, $\mathbb{M}|_{\mathcal{X}_{t}}:\mathcal{X}_{t}\to\mathcal{Y}_{t}$ is a bijection. Call $\mathcal{L}:\mathcal{Y}_{t}\to\mathcal{X}_{t}$ its (bounded) inverse and consider the equation
$\xi-\mathcal{L}(0,\xi)=\mathcal{L}(f,g).$ (22)
Here, the boundedness of $\mathcal{L}$ follows from the inverse mapping theorem [brezis2011functional, Corollary 2.7]. We see that $\xi\in\mathcal{X}_{t}$ solves (14) if and only if $\xi$ solves (22). Note that $\mathcal{L}(0,\xi)$ is a well-defined element of $\mathcal{X}_{t}$ as $(0,\xi)\in\mathcal{Y}_{t}$.
Now we apply the Fredholm alternative [gilbarg1983trudinger, Theorem 5.3] to conclude the solvability of (22). Define $T:\mathcal{X}_{t}\ni g\mapsto\mathcal{L}[0,g]\in\mathcal{X}_{t}$. Since $\mathcal{L}:\mathcal{Y}_{t}\to\mathcal{X}_{t}$ is bounded and the map
${\bf i}:\mathcal{C}^{2,\alpha}(\overline{\Omega})\cap\mathcal{X}_{t}\ni g\mapsto g\in\mathcal{C}^{1,\alpha}(\partial\Omega)\cap\left\{\xi\in\mathcal{C}^{1,\alpha}(\partial\Omega):\int_{\partial\Omega_{t}}p_{t}\xi(\nabla\phi_{t}^{*})=0\right\}$
is compact, the identity $T(g)=\mathcal{L}(0,{\bf i}(g))$ yields the compactness of $T$. Since (22) is equivalent to
$\xi-T\xi=\mathcal{L}(f,g),$ (23)
the Fredholm alternative implies that (23) admits a unique solution $\xi\in\mathcal{X}_{t}$ if and only if the homogeneous equation
$\xi-T\xi=0$ (24)
has only a trivial solution. As (24) is equivalent to (14), according to Lemma 5.10, $\xi=0$ is the unique solution of (14) in $\mathcal{X}_{t}$. The result follows. ∎
5.3 The Implicit Function Theorem
Theorem 5.6 guarantees the invertibility of $D_{\phi}\Gamma_{t}(\phi_{t})$ between the spaces $\mathcal{X}_{t}$ and $\mathcal{Y}_{t}$. However, the range of $\Gamma$ may be beyond $\mathcal{Y}_{t}$. Therefore, we modify the functional as follows to satisfy the compatibility conditions (CC). Set $t\in I$ and consider
$\tilde{\Gamma}^{(t)}:I\times(\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\cap\mathcal{X}_{t})$ $\to\mathcal{Y}_{t}$
$\left(\begin{array}[]{c}s\\ \phi\end{array}\right)$ $\mapsto\left(\begin{array}[]{c}\Gamma_{s}^{(1)}(\phi)-\int_{\Omega}q\Gamma_{s}^{(1)}(\phi)+\int_{\partial\Omega_{t}}p_{t}[\Gamma_{s}^{(2)}(\phi)](\nabla\phi_{t}^{*})\\ \Gamma_{s}^{(2)}(\phi)\end{array}\right).$ (29)
The following result shows that for each $t\in I$, the implicit function theorem (see e.g., [gilbarg1983trudinger, Theorem 17.6] or [deimling2013nonlinear, Theorem 15.1 and Corollary 15.1]) can be applied over $\tilde{\Gamma}^{(t)}$.
Lemma 5.11.
Let $t\in I$ be as above. Under the assumptions of Theorem 3.2, the following holds:
  1. (1)
    $\tilde{\Gamma}^{(t)}_{s}(\phi_{s})=0$ if and only if $\Gamma_{s}(\phi_{s})=0$.
  2. (2)
    There exist an open set $I^{\prime}\times\mathcal{U}_{t}\subset I\times(\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\cap\mathcal{X}_{t})$ with $(t,\phi_{t})\in I\times\mathcal{U}_{t}$ such that $\tilde{\Gamma}^{(t)}$ is $\mathcal{C}^{1}$ in $I^{\prime}\times\mathcal{U}_{t}$.
  3. (3)
    The Fréchet derivative of $\tilde{\Gamma}^{(t)}$ w.r.t. $\phi\in\mathcal{X}_{t}$ at $(t,\phi_{t})\in I\times(\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})\cap\mathcal{X}_{t})$, namely $D_{\phi}\tilde{\Gamma}_{t}(\phi_{t})$, satisfies $D_{\phi}\tilde{\Gamma}_{t}(\phi_{t})h=D_{\phi}\Gamma_{t}(\phi_{t})h$ for all $h\in\mathcal{X}_{t}$.
  4. (4)
    The bounded linear operator $D_{\phi}\tilde{\Gamma}_{t}(\phi_{t}):\mathcal{X}_{t}\to\mathcal{Y}_{t}$ is invertible.
Proof.
The second point is a direct consequence of Lemma 5.5 and the fourth point holds due to the third point and Theorem 5.6.
Proof of (1). We prove that if $\phi\in\mathcal{C}^{2,\alpha}_{>0}(\bar{\Omega})$ solves
$\det(D^{2}\phi)=C\frac{q}{p_{s}(\nabla\phi)}\quad{\rm in}\,\,\Omega,\ h_{s}(\nabla\phi)=0\quad{\rm in}\,\,\partial\Omega,$
for some $C\in\mathbb{R}$, then $C=1$. Note that $h_{s}(\nabla\phi)=0$ in $\partial\Omega$ implies $\nabla\phi(\Omega)=\Omega_{s}$, then
$C=\int_{\Omega}\det(D^{2}\phi)p_{s}(\nabla\phi)=\int_{\Omega_{s}}p_{s}=1$
and the claim follows.
Proof of (3). From the linearity of integration, one have for $\xi\in\mathcal{X}_{t}$
$D_{\phi}\tilde{\Gamma}^{(t)}_{t}(\phi_{t})(\xi)=\left(\begin{array}[]{c}D_{\phi}\Gamma_{t}^{(1)}(\phi_{t})(\xi)-\int_{\Omega}qD_{\phi}\Gamma_{t}^{(1)}(\phi_{t})(\xi)+\int_{\partial\Omega_{t}}p_{t}[D_{\phi}\Gamma_{t}^{(2)}(\phi_{t})(\xi)](\nabla\phi_{t}^{*})\\ D_{\phi}\Gamma_{t}^{(2)}(\phi_{t})(\xi)\end{array}\right).$
In view of $D_{\phi}\Gamma_{t}(\phi_{t})(\mathcal{X}_{t})=\mathcal{Y}_{t}$ by Theorem 5.6,
$\int_{\Omega}qD_{\phi}\Gamma_{t}^{(1)}(\phi_{t})(\xi)=\int_{\partial\Omega_{t}}p_{t}[D_{\phi}\Gamma_{t}^{(2)}(\phi_{t})(\xi)](\nabla\phi_{t}^{*}),$
then the result follows. ∎
The implicit function theorem in Banach spaces (see, e.g., [deimling2013nonlinear, Theorem 15.1]) implies that $I\ni t\mapsto\phi_{t}\in\mathcal{C}^{2,\alpha}(\overline{\Omega})$ is of class $\mathcal{C}^{1}$, meaning that $\{\phi_{t}\}_{t\in I}$ is a $\mathcal{C}^{1}$ curve. Moreover, it holds that
$D_{\phi}\tilde{\Gamma}_{t}^{(t)}(\phi_{t})](\partial_{t}\phi_{t})=-D_{t}\tilde{\Gamma}_{t}^{(t)}(\phi_{t}),$
so that $\partial_{t}\phi_{t}$ is the unique solution of
${\rm tr}([D^{2}\phi_{t}]^{-1}D^{2}\xi)+\frac{\langle\nabla p_{t}(\nabla\phi_{t}),\nabla\xi\rangle}{p_{t}(\nabla\phi_{t})}$ $=-\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}-\int_{\Omega}q\frac{\partial_{t}p_{t}(\nabla\phi_{t})}{p_{t}(\nabla\phi_{t})}+\int_{\partial\Omega_{t}}p_{t}{\partial_{t}h_{t}}\quad\rm in\,\,\Omega,$ (30)
$\langle\nabla h_{t}(\nabla\phi_{t}),\nabla\xi\rangle$ $=-\partial_{t}h_{t}(\nabla\phi_{t})\quad\rm on\,\,\partial\Omega.$
We focus now on simplifying the right-hand side of (30). To do so, we prove
$\int_{\Omega_{t}}q\partial p_{t}=\int_{\partial\Omega_{t}}p_{t}{\partial_{t}h_{t}},$ (31)
which follows as a consequence of the following Reynolds transport theorem (see e.g., [evans2022partial, Appendix C.4]) for implicitly defined surfaces.
Lemma 5.12.
Let $\{h_{t}\}_{t\in I}$, $\{\Omega_{t}\}_{t\in I}$ and $\Omega$ be as in Theorem 3.2. Then for every $\mathcal{C}^{1}$ curve $\{f_{t}\}_{t\in I}$ on $\mathcal{C}(\Omega^{\prime})$ it holds that
$\partial_{t}\int_{\Omega_{t}}f_{t}=\int_{\Omega^{\prime}}\partial_{t}f_{t}-\int_{\partial\Omega_{t}}f_{t}\partial_{t}h_{t}.$
Proof.
Set $t\in I$ and define the curve $T_{s}=(\nabla h_{t+s})^{-1}\nabla h_{t}$. By assumption and by the inverse function theorem, $\{T_{s}\}_{s\in I}\subset\mathcal{C}(\Omega^{\prime})$ is $\mathcal{C}^{1}$, $h_{t+s}(T_{s})=0$ in $\partial\Omega_{t}$, and $T_{s}(\Omega_{t})=\Omega_{s}$. Applying the second-order Taylor development on $h_{t+s}$ provides the existence of a curve $\{\omega_{s}\}_{s\in I}$ on $\mathcal{C}(\partial\Omega_{t})$ such that $o(\|\omega_{s}\|_{\mathcal{C}(\partial\Omega_{t})})=o(s)$ and
$h_{t}=0=h_{t+s}(T_{s})=h_{t+s}+\langle\nabla h_{t+s},T_{s}-T_{0}\rangle+\omega_{s}\quad\text{on}\ \partial\Omega_{t}.$
As a consequence, we get the relation
$\partial_{t}h_{t}=-\langle\left.\partial_{s}\right|_{s=0}T_{s},\nabla h_{t}\rangle\quad\text{on}\ \partial\Omega_{t}.$ (32)
Since $\nabla h_{t}$ is the unit outer-normal vector field to $\partial\Omega_{t}$, Reynolds transport theorem implies that
$\partial_{t}\int_{\Omega_{t}}f_{t}$ $=\int_{\Omega^{\prime}}\partial_{t}f_{t}+\int_{\partial\Omega_{t}}f_{t}\left\langle\left.\partial_{s}\right|_{s=0}T_{s},\nabla h_{t}\right\rangle$
and the result follows by (32). ∎
Acknowledgments
The authors thank Eustasio del Barrio, Nicolás García Trillos, Alejandro Garriz Molina, Tudor Manole, Gilles Mordant and Marcel Nutz for helpful comments.